### v2

# Better bootstrap $t$ confidence intervals for the mean

Abstract
This article explores combinations of weighted bootstraps, like the Bayesian bootstrap, with the bootstrap $t$ method for setting approximate confidence intervals for the mean of a random variable in small samples. For this problem the usual bootstrap $t$ has good coverage but provides intervals with long and highly variable lengths. Those intervals can have infinite length not just for tiny $n$, when the data have a discrete distribution. The BCa bootstrap produces shorter intervals but tends to severely under-cover the mean. Bootstrapping the studentized mean with weights from a Beta(1/2,3/2)(1/2,3/2) distribution is shown to attain second order accuracy. It never yields infinite length intervals and the mean square bootstrap $t$ statistic is finite when there are at least three distinct values in the data, or two distinct values appearing at least three times each. In a range of small sample settings, the beta bootstrap $t$ intervals have closer to nominal coverage than the BCa and shorter length than the multinomial bootstrap $t$. The paper includes a lengthy discussion of the difficulties in constructing a utility function to evaluate nonparametric approximate confidence intervals.
1 Introduction
This article is about nonparametric approximate confidence intervals (ACIs) for the mean of a real random variable. There are two leading bootstrap solutions for this problem. One is the bootstrap $t$ method and the other is the bias corrected accelerated (BCa) method. These are both well understood theoretically in the limit as the sample size $n$ of an IID sample diverges to infinity. Under mild conditions on the sampling distribution, both are second order accurate with both one and two-sided coverage errors decaying as $O(1/n)$. Alternative methods that are not second order accurate typically have one-sided coverage errors of $O(1/\sqrt{n})$ with two-sided errors of $O(1/n)$ (Hall,, 1988).
For smaller values of $n$ there is a nasty tradeoff. The bootstrap $t$ intervals come much closer to attaining the desired nominal coverage, but they do so at the expense of being quite long. The BCa intervals are narrower than the bootstrap $t$ intervals, but they very often have coverage levels severely below the nominal target. This article develops some alternative bootstrap $t$ confidence intervals that retain the favorable asymptotic coverage while being noticeably narrower than the customary bootstrap $t$ intervals in a collection of small sample settings. These methods attain much higher coverage than the BCa intervals.
One explanation for the wide intervals from the bootstrap $t$ method is that (as described below) some resamplings place zero weight on some of the $n$ observations. When one or more extreme observations gets zero weight, the resampled standard deviation becomes small which can make the sample mean appear to be an outlier. We then get very wide tails for the resampled $t$ distribution, including atoms at $\pm\infty$. The result is wide confidence intervals for the mean and consequently, high coverage.
We can eliminate the zero weights by merging the Bayesian bootstrap (Rubin,, 1981) with the bootstrap $t$. That provides asymptotically valid confidence intervals, though it lacks second order accuracy for one-sided confidence intervals. Unfortunately, the Bayesian bootstrap version of the bootstrap $t$ intervals does not have very good small sample coverage in the cases examined here. Perhaps the good coverage properties of the original bootstrap $t$ are strongly dependent on those small weights. A next step is to then to replace the unit exponential weights of the Bayesian bootstrap by weights from a distribution that has a singular probability density at the origin in order to get near zero weights that remain positive. Strictly positive weights keep the resampled variance positive unless all $n$ observations are equal. That positive resample variance then keeps the resampled $t$ value finite. One such weight distribution is Beta($1/2,3/2$). We show that it has a second order accuracy property as $n\to\infty$, so its small sample advantages do not conflict with asymptotic performance.
If having some near zero weights works well, we might expect that having many near zero weights could work even better. One way to do that is to use a very heavy-tailed weight distribution such that the largest sample value is often large compared to the sum of the other n1n-1 values. That is not viable in the case of bootstrap weights. As noted below the weights must have finite variance prior to normalization and that precludes the largest one from dominating the others, at least for large $n$. We will also see a reason to have a finite fourth moment for the weights. A very heavy-tailed weight distribution, the lognormal, is seen to give very short bootstrap $t$ confidence intervals, often shorter than even the BCa intervals, and with lower coverage.
It is also possible to resample with even more zero weights than the multinomial distribution provides. The Poisson distribution with mean $1$ has been used for bootstrap weights in Oza and Russell, (2001). While its asymptotic probability of a zero weight matches that of multinomial sampling, it has higher probability of a zero weight for the small $n$ that we consider here. Hartigan, (1969) considers sampling from all $2^{n}$ subsets of the data which we can implement by giving each observation weight 0 or $1$ with probability $1/2$ each. To keep the mean weight equal to one, we can equivalently give weight 0 or $2$. Owen and Eckles, (2012) refer to this as the double-or-nothing bootstrap. Here we use it in a bootstrap $t$ along with a half-sampling bootstrap $t$ described below. These methods with more exact zero weights give even wider confidence intervals than the multinomial bootstrap $t$ does. As a result, we can get a spectrum of bootstrap $t$ intervals ranging from some that are even narrower than the BCa intervals to others that are wider than the usual multinomial bootstrap $t$.
For continuous data, it is easy to see that the probability of an infinite resampled $t$ statistic is $n^{1-n}$ which rapidly becomes negligible compared to a threshold like $0.025$ as $n$ increases. What is less remarked on is that this probability can decay much more slowly for discrete data. In some computations below we see infinite length intervals for the usual bootstrap $t$ with $n=19$ for the mean of $\mathrm{Pois}(1)$ data. With $\mathrm{Beta}(1/2,3/2)$ weights, the bootstrap $t$ statistic is finite. It can have heavy tails for small $n$. However once there are three or more distinct observations, or two distinct observations seen at least three time each, the expected squared bootstrap $t$ value given the data, becomes finite.
Most of what is known about bootstrap confidence intervals comes from asymptotic expansions as $n\to\infty$, using assumptions about the moments of the sampling distribution. Even for just one given bootstrap method there can be significant differences in small sample performance for different distributions having the same low order moments. An extreme version of this arises in the behavior of the bootstrap for discrete data. A Cramér condition that
$\limsup_{|t|\to\infty}|\mathbb{E}(\exp(itx))|<1$ (1)
plays a large role in the asymptotic theory of Hall, (1986). The assumption provides more tractable Edgeworth expansions. Equation (1) rules out cases where the support of $x$ is confined to an arithmetic sequence, so it is violated for integer-valued random variables such as counts.
We are left with an unsatisfactory situation. The cases with small $n$ have the most consequential differences and the least theoretical understanding. The study for small values of $n$ has primarily been done numerically. Numerical investigation still seems to be necessary, nearly fifty years after the bootstrap appeared.
As mentioned above, strictly positive weights keep the interval lengths finite. Some other advantages of strictly positive reweighting arise in other contexts. Xu et al., (2020) mention several. When every original data point is at least partially represented in every bootstrap sample, then design matrices for regression that are non-singular remain nonsingular in every bootstrap sample. If a binary response is not linearly separable as a function of the predictors in the sample, then this continues to hold true in bootstrap samples where every observation has positive weight. They also mention similar advantages for survival analysis problems where the majority of the lifetimes have been censored. Mase et al., (2024) use positive weights for a variable importance problem so that while bootstrapping the importance of variable $j$ to an outcome for subject $i$, that subject $i$ is represented in every bootstrap sample. Owen, (2009) finds that quasi-Monte Carlo sampling with the Bayesian bootstrap weights is very effective for bootstrap bias estimation.
A great difficulty in comparing nonparametric ACIs for the mean is that there is no universally accepted figure of merit with which to compare them. It is impossible to get exactly the nominal coverage and so we must choose between methods with higher coverage and methods with shorter intervals. There are many rules of thumb and default choices in statistics but none that say how much wider an interval could reasonably be in return for reducing undercoverage by 1%. Without a quantitative answer we can only look at Pareto frontiers and make subjective judgments.
An outline of this paper is as follows. Section 2 sets up the notation, describes BCa intervals and weighted bootstrap $t$ intervals and gives some choices for the weights. Section 3 shows that a bootstrap $t$ using Beta($1/2,3/2$) weights meets the criteria for second order accuracy in Hall and Mammen, (1994). It also has a finite expected mean squared bootstrap $t$ statistic given data with three or more distinct values, or two distinct values appearing at least three times each, and it compares interval lengths for the case with $n=2$ distinct observations. Section 4 describes in detail the difficulties in defining a common task framework to judge ACIs, considering calibration, elicitation, decision theory and a similar tradeoff that arises for precision and recall in classification. Section 5 presents a collection of test cases for investigations of bootstrap ACIs with numerical comparisons for sample sizes $2\leqslant n\leqslant 20$. It makes a graphical comparison of the methods most of which fall on a Pareto frontier trading off interval length and coverage. Section 6 presents some conclusions. Appendix A contains some proofs. Appendix B surveys some of the literature on the bootstrap for small $n$. Appendix C describes conventions for handling some corner cases that only have a meaningfully large probability when $n$ is as small as we consider here.
2 Background and notation
We suppose throughout that $x_{1},\dots,x_{n}$ are sampled IID from a distribution $F$ with mean $\mu$ and variance $\sigma^{2}\in(0,\infty)$. A two-sided confidence interval for $\mu$, at level $\alpha\in(0,1)$, is a pair of random variables $L$ and $U$, given as functions of $x_{1},\dots,x_{n}$ for which $\Pr(L\leqslant\mu\leqslant U)=\alpha$. We always have $L\leqslant U$. A customary default choice is $\alpha=0.95$. While $0.95$ is not a very high confidence level, it is already a challenge to get close to this coverage in small samples, even for two-sided coverage.
The two-sided intervals we consider are formed from a pair of one-sided intervals. We select $0<\alpha_{1}<\alpha_{2}<1$ and seek random $L$ and $U$ with $\Pr(\mu<L)=\alpha_{1}$ and $\Pr(\mu\leqslant U)=\alpha_{2}$ so that $\Pr(L\leqslant\mu\leqslant U)=\alpha_{2}-\alpha_{1}$. For equal one-sided coverage errors we take $\alpha_{1}=(1-\alpha)/2$ for $\alpha>1/2$ and $\alpha_{2}=1-\alpha_{1}=(1+\alpha)/2$. Then the customary values are $\alpha_{1}=0.025$ and $\alpha_{2}=0.975$.
Although we want an exact confidence interval for $\mu$, we can only have one under special circumstances. Confidence intervals are possible within parametric families of distributions, but then in addition to $1-\alpha$ uncertainty from the interval, we have uncertainty about the extent to which model violations affect our interval. In a Bayesian setting we may well obtain a credible interval with our desired coverage level. That calibration extends to others who share the same prior distribution, but not universally. Bayesian methods are often used pragmatically with noninformative or weakly informative priors in order to improve upon frequentist methods. For example, Greenland and Poole, (2013) note that under weak priors “95% confidence intervals approximate 95% posterior probability intervals”. Such intervals are also not exact and we face the same tradeoff of calibration versus length.
While we prefer a nonparametric confidence interval, those are essentially ruled out by a theorem of Bahadur and Savage, (1956). Here is the description from Owen, (2024):
They consider a set $\mathcal{F}$ of distributions on $\mathbb{R}$. Letting $\mu(F)$ be $\mathbb{E}(Y)$ when $Y\sim F\in\mathbb{R}$, their conditions are:
  1.  (i)
    For all $F\in\mathcal{F}$, $\mu(F)$ exists and is finite.
  2.  (ii)
    For all $m\in\mathbb{R}$ there is $F\in\mathcal{F}$ with $\mu(F)=m$.
  3.  (iii)
    $\mathcal{F}$ is convex: if $F,G\in\mathcal{F}$ and $0<\pi<1$, then $\pi F+(1-\pi)G\in\mathcal{F}$.
Then their Corollary 2 shows that a Borel set constructed based on $Y_{1},\dots,Y_{N}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$ that contains $\mu(F)$ with probability at least $1-\alpha$ also contains any other $m\in\mathbb{R}$ with probability at least $1-\alpha$. More precisely: we can get a confidence set, but not a useful one. They allow $N$ to be random so long as $\Pr(N<\infty)=1$.
In special settings, we know that $0\leqslant x\leqslant 1$ and in those cases it is possible to construct confidence intervals using Hoeffding’s theorem or using empirical Bernstein intervals. See Waudby-Smith and Ramdas, (2024) for a survey and some new methods that are asymptotically narrower than predecessors. While these methods attain at least the desired coverage, they are much wider than the intervals based on the CLT (Austern and Mackey,, 2022). Similarly, if $x$ has a symmetric distribution with $\mathbb{E}(|x|)<\infty$ then $\mathbb{E}(x)$ equals the median of $x$ and it is possible to get a confidence interval for it.
When $F$ does not belong to a known parametric family and $x$ does not satisfy some special side condition (such as known bounds or symmetry) then we must rely instead on approximate confidence intervals (ACIs). These are usually based on the central limit theorem where $\sqrt{n}(\bar{x}-\mu)/\sigma\stackrel{{\scriptstyle\cdot}}{{\sim}}\mathcal{N}(0,1)$ for ‘large enough’ $n$. There is no sharp distinction where we can say that $n$ is large enough while n1n-1 isn’t. A very common rule of thumb is to consider $n\geqslant 30$ large enough. Kufs, (2010) looks into the history of this rule of thumb. Miller, (1986) even suggests that $n\geqslant 10$ may be enough, if one can accept mild skewness in $\bar{x}$. On the other hand Hesterberg, (2015) has stringent conditions (one-sided coverages within $0.0025$ of $0.025$) where Student’s $t$ needs $n\geqslant 4815$ for $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Exp}(1)$. Based on some simulations, Boos and Hughes-Oliver, (2000) suggest $n\geqslant(5.66\gamma)^{2}$ to get nominal 95% intervals Student’s $t$ intervals with two-sided coverage at least 9494%, when $x_{i}$ have skewness $\gamma$. Kohavi et al., (2014) find that for some highly skewed electronic commerce data this guideline requires $n\geqslant 114{,}000$. These higher numbers provide good motivation for further investigation of confidence intervals.
The suitability of these rules of thumb depends on how non-Gaussian the distribution of $x$ is and how harmful it is to have less than nominal coverage. In practice we must choose an ACI without precise information about these.
We will look at performance of bootstrap ACIs for $n$ in the range from $2$ to $20$. It is not reasonable to expect good results for $n=2$. Nor is there a sharp line where we can say that a reasonable result can be obtained for $n\geqslant n_{0}$ but not for $n\leqslant n_{0}-1$. The bootstrap methods we consider will all perform very well as $n\to\infty$ and will all be problematic for $n=2$ on difficult distributions. What we can do is observe the trajectory of their length and coverage as a function of $n$ and see which of them encounter less difficulty when $n$ is small. Investigators may have some idea which example settings are most similar to their problem, even if that knowledge cannot be expressed precisely.
2.1 Approximate confidence intervals
The ACIs we consider are based on the sample mean and variance
$\bar{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}\quad\text{and}\quad\hat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2},$
and bootstrap versions of these quantities. The unbiased sample variance estimate is $s^{2}=n\hat{\sigma}^{2}/(n-1)$. It figures in the sample $t$ quantity
$t=\sqrt{n}\frac{\bar{x}-\mu}{s}$
in which $\mu$ is a hypothesized value for $\mathbb{E}(x)$.
The best known ACI for the mean has upper and lower limits
$\bar{x}\pm st^{1-\alpha/2}_{(n-1)}/\sqrt{n}$
where $t^{1-\alpha/2}_{(n-1)}$ is the $1-\alpha/2$ quantile of Student’s $t$ distribution on n1n-1 degrees of freedom. It satisfies
$\lim_{n\to\infty}\Pr\Bigl{(}\bar{x}-\frac{s}{\sqrt{n}}t^{1-\alpha/2}_{(n-1)}\leqslant\mu\leqslant\bar{x}+\frac{s}{\sqrt{n}}t^{1-\alpha/2}_{(n-1)}\Bigr{)}=1-\alpha$
by the central limit theorem and it is exact when $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(\mu,\sigma^{2})$ and $n\geqslant 2$.
For small samples from a distribution $F$ that includes atoms we must allow for the possibility that $\hat{\sigma}=0$. We adopt a convention that the resulting ACI is simply the point $\bar{x},\bar{x}]$. Any value of $\mu$ other than $\bar{x}$ would have given $|t|=\infty$. This is the first of several conventions that we need to use in order to handle very small $n$. They are collected in Appendix C.
2.2 Bootstrap $t$
Here we present the bootstrap $t$ algorithm of Efron, (1981). It is also known as the percentile $t$ algorithm (Hall,, 1986).
For our purposes of generalizing the bootstrap $t$, it is more convenient to use $\hat{\sigma}^{2}$ than $s^{2}$. The corresponding $t$ statistic is
$\tilde{t}=\sqrt{n}\frac{\bar{x}-\mu}{\hat{\sigma}}=\sqrt{\frac{n}{n-1}}t.$
To get a bootstrap sample $x_{1}^{*},\dots,x_{n}^{*}$ we draw indices $j(1),\dots,j(n)$ IID from $\mathbb{U}\{1,2,\dots,n\}$ and set $x_{i}^{*}=x_{j(i)}$ for $i=1,\dots,n$. This amounts to resampling the original $x_{i}$ values with replacement. From the bootstrap data we can compute
$\bar{x}^{*}=\frac{1}{n}\sum_{i=1}^{n}x_{i}^{*},\quad\hat{\sigma}^{*2}=\frac{1}{n}\sum_{i=1}^{n}(x_{i}^{*}-\bar{x}^{*})^{2}\quad\text{and}\quad t^{*}=\sqrt{n}\frac{\bar{x}^{*}-\bar{x}}{\hat{\sigma}^{*}}.$ (2)
The bootstrap $t$ ACI is based on equating the unknown distribution of $\tilde{t}=\sqrt{n}(\bar{x}-\mu)/\hat{\sigma}$ under $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$ to the distribution of $t^{*}$ from equation (2). We write $t^{*}$ instead of $\tilde{t}^{*}$ since the presence of the asterisk already indicates that it is a bootstrap quantity. The distribution of $t^{*}$ can be computed to any desired level of accuracy by sampling data sets $(x_{1}^{*},\dots,x_{n}^{*})$ some number $B$ of times. A common choice is $B=2000$. The resulting ACI is
$\bigl{[}\bar{x}-\hat{\sigma}t^{(*\alpha_{2})}/\sqrt{n},\bar{x}-\hat{\sigma}t^{(*\alpha_{1})}/\sqrt{n}\,\bigr{]}$ (3)
where $t^{(*\alpha_{j})}$ is the $\alpha_{j}$ quantile of the distribution of $t^{*}$. It might be more precise to write $\hat{t}^{(*\alpha_{j})}$, as it is an estimated quantile but we will write $t^{(*\alpha_{j})}$ to get less cluttered notation.
Hall and Mammen, (1994) consider several ways to define a bootstrap $t$. When their weights $N_{ni}^{*}$ are normalized to sum to $n$, like $nw_{i}$ here, then their $\hat{\sigma}^{*2}$ on page 2017 reduces to $\hat{\sigma}^{*2}$ above. They refer $(\bar{x}-\mu)/\hat{\sigma}$ to $(\bar{x}^{*}-\bar{x})/\hat{\sigma}^{*}$ which is the same bootstrap approximation we use here. Mason and Newton, (1992) study a bootstrap $t$ method where $\hat{\sigma}^{*}$ is multiplied by a standard deviation of the resampling weights. The multinomial bootstrap $t$ is not then a special case of their quantity. Csörgő et al., (2014) consider this same statistic as well as $\sqrt{\sum_{i=1}^{n}v_{i}}(\bar{x}^{*}-\bar{x})/\hat{\sigma}^{*}$. That latter statistic does generalize $t^{*}$. They show consistency but do not consider convergence rates.
The bootstrap $t$ is based on a plug-in principle. If we knew the distribution of $(\bar{x}-\mu)/\hat{\sigma}$ for $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$, we could use it along with the known values of $\bar{x}$ and $\hat{\sigma}$ to get exact confidence intervals for $\mu$. We don’t know that distribution but we can get the distribution of $(\bar{x}^{*}-\bar{x})/\hat{\sigma}^{*}$ for $x_{i}^{*}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\widehat{F}$ where $\widehat{F}$ is the empirical distribution of $x_{1},\dots,x_{n}$. So we ‘plug-in’ $\widehat{F}$ for $F$ to approximate the distribution of $(\bar{x}-\mu)/\hat{\sigma}$ by that of $(\bar{x}^{*}-\bar{x})/\hat{\sigma}^{*}$ and this gives us an ACI for $\mu$.
There is positive probability that $\hat{\sigma}^{*}=0$. This will ordinarily give $t^{*}=\pm\infty$, resulting in wider ACIs. With no ties in the data, $\Pr(\hat{\sigma}^{*}=0)=n^{1-n}$ which for moderately large $n$ is typically much smaller than $\min(\alpha_{1},1-\alpha_{2})$. Then retaining some infinite values of $|t^{*}|$ still results in a finite ACI length that is wider than it would be if we had omitted the infinite cases. Some weighted bootstraps that we consider below have larger values of $\Pr(\hat{\sigma}^{*}=0)$ than the usual bootstrap $t$ has, and we see wider ACIs for them. For discrete data, such as counts, the possibility of $\hat{\sigma}^{*}=0$ is greatly increased.
When there are ties in the data, it is possible to have $\bar{x}^{*}-\bar{x}=\hat{\sigma}^{*}=0$. We adopt a convention that $t^{*}=0$ in this case. For a method with more frequent occurences of $\bar{x}^{*}-\bar{x}=\hat{\sigma}^{*}=0$, the distribution of $t^{*}$ will have a larger atom at 0. When $\Pr(x=\mu)=0$, this will result in shorter ACIs with reduced coverage of the true mean, leading us to prefer a different method.
Instead of the bootstrap $t$, one could simply use the quantiles of $\bar{x}^{(*\alpha_{j})}$ themselves to form an ACI for $\mu$. This is known as the percentile method. It is not competitive with the bootstrap $t$ method when we want an ACI for the mean. It is not as well regarded as the BCa method described next. The percentile method and some other bootstraps are described in Efron and Tibshirani, (1993).
2.3 Bias corrected accelerated bootstrap
The bias corrected accelerated, or BCa, ACI is from Efron, (1987). Here we consider the version used in the R function bcanon from the bootstrap package Leisch, (2019) based on S functions by R. Tibshirani for the book Efron and Tibshirani, (1993). We describe it for a sample mean though it can be applied to more general functions.
The BCa interval uses the same values $x_{1}^{*},\dots,x_{n}^{*}$ and their average $\bar{x}^{*}$ that the bootstrap $t$ does. Given $B$ values of $\bar{x}^{*}$ it sets a bias adjustment parameter
$z_{0}=\Phi^{-1}\Bigl{(}\frac{1}{B}\sum_{b=1}^{B}1\{\bar{x}^{*b}<\bar{x}\}\Bigr{)}.$ (4)
Then it uses leave one out means $\bar{x}_{-i}=(n\bar{x}-x_{i})/(n-1)$ to define an acceleration quantity
$a=\frac{\sum_{i=1}^{n}u_{i}^{3}}{6\bigl{(}\sum_{i=1}^{n}u_{i}^{2}\bigr{)}^{3/2}}$
for $u_{i}=\bar{x}-\bar{x}_{-i}=(x_{i}-\bar{x})/(n-1)$. Then modifying the limits $\alpha_{j}$ for $j=1,2$ to
$\tilde{\alpha}_{j}=\Phi\Bigl{(}z_{0}+\frac{z_{0}+z_{j}}{1-a(z_{0}+z_{j})}\Bigr{)}$
the lower and upper confidence limits are the quantiles $\bar{x}^{*(\tilde{\alpha}_{j})}$ for $j=1,2$.
The BCa method is harder to interpret than the bootstrap $t$. It is derived using the idea that there exists a monotone increasing transformation $\phi(\cdot)$ for which $\phi(\bar{x})\sim\mathcal{N}(\phi(\mu)-z_{0}\sigma_{\phi},\sigma^{2}_{\phi})$ where $\sigma_{\phi}=1+a\phi(\mu)$. The form of the acceleration is based on an asymptotic argument to adjust the confidence level to be second order accurate. DiCiccio and Efron, (1996) remark that the BCa intervals strike some people as being ‘unbootstraplike’ because the formula for $a$ does not come from the bootstrap replications.
To use the BCa interval we only need to be able to calculate our scalar statistic of interest on bootstrap samples of the data. The bootstrap $t$ requires a quantity to use for $\hat{\sigma}$. When the statistic of interest is not a sample mean, then the BCa intervals remain easy to use while the bootstrap $t$ can become difficult to use.
2.4 Bootstrap as reweighting
In the bootstrap sample $x_{1}^{*},\dots,x_{n}^{*}$ the observation $x_{i}$ is included a random number $v_{i}$ of times. Then $\bar{x}^{*}=\sum_{i=1}^{n}w_{i}x_{i}$ where the weights $w_{i}$ are normalized counts: $w_{i}=v_{i}/\sum_{i^{\prime}=1}^{n}v_{i^{\prime}}=v_{i}/n$. The distribution of each $v_{i}$ is $\mathrm{Bin}(n,1/n)$ and the joint distribution of $(v_{1},\dots,v_{n})$ is a multinomial.
The Bayesian bootstrap of Rubin, (1981) uses different weights. It samples $v_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Exp}(1)$ and then takes $w_{i}=v_{i}/\sum_{i^{\prime}=1}^{n}v_{i^{\prime}}$. The pseudo-counts $v_{i}$ are positive non-integer values. This bootstrap has a connection to Bayesian statistics and is usually used in percentile methods.
We can write the bootstrap $t$ method in terms of weighted resampling for general bootstrap weights. For that, we take
$\hat{\mu}^{*}=\sum_{i=1}^{n}w_{i}x_{i},\quad\hat{\sigma}^{2*}=\sum_{i=1}^{n}w_{i}(x_{i}-\hat{\mu}^{*})^{2}\quad\text{and}\quad t^{*}=\sqrt{n}\frac{\hat{\mu}^{*}-\bar{x}}{\hat{\sigma}^{*}}.$
Then the ACI is
$\Bigl{[}\bar{x}-\frac{\hat{\sigma}}{\sqrt{n}}t^{(*\alpha_{2})},\bar{x}-\frac{\hat{\sigma}}{\sqrt{n}}t^{(*\alpha_{1})}\Bigr{]}.$ (5)
The quantiles of $t^{*}$ are estimated by sampling pseudo-counts, normalizing them into weights and then computing the weighted mean, variance and $t$ statistic. The bootstrap $t$ ACIs we consider take the form (5) for different distributions of $(v_{1},\dots,v_{n})$.
In the weighting approach we are not exactly mimicking the sampling process that lead to the original data $x_{i}$ by plugging in $\widehat{F}$ for $F$. We are however retaining the idea of randomly reweighting the data that the original bootstrap has. In that sense, it is bootstraplike.
We can use some other weights than the multinomial or exponential ones but the weights cannot be completely arbitrary. For properly calibrated asymptotic coverage it suffices for IID pseudocounts $v_{i}$ to have mean $1$ and variance $1$. For second order accuracy we also want skewness one (Hall and Mammen,, 1994). That is, we want $\mathbb{E}(v_{i})=\mathbb{E}((v_{i}-1)^{2})=\mathbb{E}((v_{i}-1)^{3})=1$. We will only consider $v_{i}\geqslant 0$, because then $\Pr(\hat{\sigma}^{*2}<0)=0$.
The assumption that $\mathbb{E}(v_{i})=\mathrm{var}(v_{i})=1$ has a consequence in terms of effective sample size. Conditionally on the values of pseudo-counts $v_{i}$, the variance of the weighted mean $\sum_{i=1}^{n}v_{i}x_{i}/\sum_{i=1}^{n}v_{i}$ is $\sigma^{2}\sum_{i=1}^{n}v_{i}^{2}/(\sum_{i=1}^{n}v_{i})^{2}$. Equating that to $\sigma^{2}/n_{\mathrm{eff}}$ gives an effective sample size of
$n_{\mathrm{eff}}=\frac{(\sum_{i=1}^{n}v_{i})^{2}}{\sum_{i=1}^{n}v_{i}^{2}}.$
We find that $n_{\mathrm{eff}}/n\to 1/2$ as $n\to\infty$ when $v_{i}$ are IID with mean $1$ and variance $1$. Thus we may think of the bootstrap as a reweighting the data using an effective sample size of about $n/2$.
The Poisson distribution with mean $1$ has variance and skewness also equal to $1$. Oza and Russell, (2001) use $v_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Pois}(1)$ in online bootstrap aggregation for machine learning. For large $n$, the multinomial weights of the bootstrap are almost equivalent to using independent Poisson weights with mean $1$. Independent Poisson weights are more convenient than multinomial weights in a streaming environment.
The random variables $v_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathbb{U}\{0,2\}$ have mean $1$ and variance $1$ and they can be used in a bootstrap. For even $n$, we can similarly select $n/2$ of the $v_{i}$ to be $2$ in a simple random sample and let the other $v_{i}$ be 0. Neither of these will lead to a second order accurate bootstrap. These methods can be derived using the typical value theory of Hartigan, (1969). A set of N1N-1 distinct ‘typical’ values partitions $\mathbb{R}$ into $N$ intervals (two of which have infinite length) such that each of the $N$ intervals has probability $1/N$ of containing the parameter of interest. For a symmetric distribution we can take means of $N=2^{n}-1$ different subsamples of the data and then form a confidence interval from the union of a ‘central’ subset of those intervals.
For large $n$, normalized $\mathbb{U}\{0,2\}$ weights correspond very closely to the half-sampling method used in survey sampling (McCarthy,, 1969) which takes simple random samples of half of the data points. In our bootstrap $t$ version of half-sampling, when $n$ is even we select $n/2$ of $n$ observations by a simple random sample where all $n\choose n/2$} selections being equally probable. When $n$ is odd we take a simple random sample of either $(n+1)/2$ points or (n1)/2(n-1)/2 points, each with probability $1/2$.
The Poisson and double-or-nothing weights can both give $v_{1}=v_{2}=\cdots=v_{n}=0$ in which case the weights $w_{i}$ are not well defined. For double-or-nothing this happens with probability $2^{-n}$ while for the Poisson weights it happens with probability $\exp(-n)$. The chance of this happening is miniscule for modestly large $n$ but for smaller $n$ this probability is not negligible compared to customary values of $\alpha_{1}$ and $1-\alpha_{2}$. A reasonable convention for this case is to discard such weights and then sample $(v_{1},\dots,v_{n})$ until $B$ cases with $\max_{i}v_{i}>0$ have been obtained.
In his work on the wild bootstrap, Mammen, (1993) presents a distribution on just two support points that has mean zero, variance one and skewness one. Adding one to this random variable gives a positive random pseudo-count $v$ that satisfies the conditions for second order accuracy. Specifically, if
$\Pr\Bigl{(}v=\frac{3\pm\sqrt{5}}{2}\Bigr{)}=\frac{1}{2}\mp\frac{1}{2\sqrt{5}},$ (6)
then $\mathbb{E}(v)=\mathrm{var}(v)=\mathbb{E}((v-1)^{3})=1$.
We can combine $\mathbb{E}(v)=\mathrm{var}(v)=\mathbb{E}((v-1)^{3})=1$ with $\Pr(v=0)>0$ by taking $v=0,1,3$ with probabilities $1/3,1/2,1/6$ respectively. This $v$ has kurtosis 0. Another solution takes $v=0,1,2,4$ with probabilities $9,8,6,1$ divided by $24$. It has kurtosis $1$ like the Poisson distribution. These combine second order accuracy with an atom at zero. We don’t consider them further because Poisson weights already have both of those properties.
There has not been much attention paid to the kurtosis of the pseudo-counts. One exception is Guillou, (1999) who gives a kurtosis condition that makes the one-sided coverage errors $O(n^{-3/2})$ for studentized statistics like we study here. In other words, she can obtain third order accuracy. In addition to a Cramer condition, Guillou requires the sample skewness of $v_{i}$ to be $1+O_{p}(n^{-1/2})$ and also requires the sample kurtosis of the $v_{i}$ to satisfy a relation with the sample kurtosis of the $x_{i}$.
2.5 Weights that are singular at the origin
We would like a distribution for pseudocounts $v$ with $\Pr(v>0)=1$ but for which $v$ takes values very close to zero with high probability, by having a probability density function that diverges to infinity as $v\to 0$. The $\mathrm{Beta}(a,b)$ distributions with first parameter $0<a<1$ are singular at the origin, and scaling them to have mean and variance equal to one, generates a parametric family of distributions indexed by $a$.
Proposition 1.
For $0<a<1$, and $b=(a^{2}+a)/(1-a)$ let $x\sim\mathrm{Beta}(a,b)$ and $v=Ax$ for A=2/(1a)A=2/(1-a). Then $\mathbb{E}(v)=\mathrm{var}(v)=1$ and $v$ has skewness $2a$.
Proof.
First a+b=2a/(1a)=aAa+b=2a/(1-a)=aA so
$\mathbb{E}(x)=\frac{a}{a+b}=\frac{1}{A}$
and then $\mathbb{E}(v)=1$. Next a+b+1=aA+1=(12/A)A+1=A1a+b+1=aA+1=(1-2/A)A+1=A-1, and so
$\mathrm{var}(x)=\frac{1}{A}\Bigl{(}1-\frac{1}{A}\Bigr{)}(a+b+1)^{-1}=\frac{1}{A}\Bigl{(}1-\frac{1}{A}\Bigr{)}\frac{1}{A-1}=\frac{1}{A^{2}}$
so $\mathrm{var}(v)=1$. Finally, $v$ and $x$ both have skewness
$\gamma(a)=\frac{2(b-a)\sqrt{a+b+1}}{(a+b+2)\sqrt{ab}}$
Now a+b+1=(1+a)/(1a)a+b+1=(1+a)/(1-a), ba=2a/(1a)b-a=2a/(1-a) and $ab=a^{2}(a+1)/(1-a)=a^{2}(a+b+1)$, so
$\gamma(a)=\frac{2(b-a)}{(a+b+2)a}=\frac{4a(1-a)^{-1}}{2(1-a)^{-1}a}=2a$
as required. ∎
There are a few convenient special cases. First, taking $a=1/2$ we get $v=4\times\mathrm{Beta}(1/2,3/2)$ with mean, variance and skewness all equal to $1$. Theorem 1 below shows that this distribution satisfies the conditions of Hall and Mammen, (1994) for second order accuracy. We call the resulting method the beta bootstrap $t$.
Another convenient choice arises with $b=1$. We get this by taking $a=\sqrt{2}-1\doteq 0.41$ and $A=2+\sqrt{2}$. Then the probability density of $v$ is
$f(v)=\frac{a}{A^{a}}v^{a-1}1_{0<v<A}.$ (7)
We call the resulting method the power bootstrap $t$ as $f$ is simply a negative power of $v$, scaled and truncated to $(0,A)$. Properties of these beta distributed weights and some other weights are summarized in Table 1.
Smaller values of $a$ correspond to beta distributions with even stronger singularities at the origin. We see below that generally results in longer ACIs and higher coverage. For instance the power bootstrap $t$ typically has longer intervals than the beta bootstrap $t$. If we take the limit as $a\to 0$ then $A\times\mathrm{Beta}(a,b)$ converges in distribution to $\mathbb{U}\{0,2\}$, so the double-or-nothing bootstrap arises as the limit of this beta family. Some other interesting examples are $3\times\mathrm{Beta}(1/3,2/3)$ and $6\times\mathrm{Beta}(2/3,10/3)$ because they have simple expressions with round numbers. When $a<\sqrt{2}-1$, we get $b<1$ and then the density is no longer monotonic decreasing in $v$. For $a\geqslant 1$, it is impossible to scale the beta distribution to have mean and variance both equal to one as this would give $b\not\in(0,\infty)$. In a mixture model with random $a\in(0,1)$ the variable $v\sim A(a)\mathrm{Beta}(a,b(a))$, has mean variance and skewness equal to $1$ if $\mathbb{E}(a)=1/2$.
Some gamma distributions have unbounded probability density functions at the origin. However the only gamma distribution with mean one and variance one is the exponential distribution used already in the Bayesian bootstrap. It has a bounded density and also has skewness $2$.
Method Distribution IID $\mathbb{E}(v)$ $\mathrm{var}(v)$ $\mathrm{skew}(v)$ $\Pr(v=0)$
Original Multinomial 1 11/n1-1/n $\frac{1-2/n}{\sqrt{1-1/n}}$ $(1-1/n)^{n}$
Half sampling Simple random 1 1 0 $1/2$
Bayesian $\mathrm{Exp}(1)$ $\checkmark$ 1 1 2 0
Poisson $\mathrm{Pois}(1)$ $\checkmark$ 1 1 1 $e^{-1}$
Beta $4\times\mathrm{Beta}(1/2,3/2)$ $\checkmark$ 1 1 1 0
Power Equation (7) $\checkmark$ 1 1 $2(\sqrt{2}-1)$ 0
Mammen Equation (6) $\checkmark$ 1 1 1 0
Double/nothing $\mathbb{U}\{0,2\}$ $\checkmark$ 1 1 0 $1/2$
Lognormal $\exp(\mathcal{N}(-\log(2)/2,\log(2)))$ $\checkmark$ 1 1 4 0
Table 1: Properties of the individual (pseudo-)counts $v_{i}$ in different bootstrap weighing schemes.
2.6 Heavy-tailed weights
A useful example of a heavy-tailed distribution is the log-normal distribution. By taking $v_{i}=\exp(-\log(2)/2+\sqrt{\log(2)}Z)$ for $Z\sim\mathcal{N}(0,1)$ we get $\mathbb{E}(v_{i})=\mathrm{var}(v_{i})=1$. Bootstrap $t$ ACIs with $v_{i}$ from this log-normal distribution will not be second order accurate because the skewness of $v_{i}$ is not equal to one. There does not seem to be a common distribution of positive $v$ with very heavy tails and mean, variance and skewness all equal to one. A bootstrap $t$ with such a distribution would be second order accurate. We will see that log-normal weights lead to very short ACIs and low coverage.
It is reasonable to expect low coverage for other heavy-tailed distributions too. One other heavy-tailed sampler takes $\Pr(v=\sqrt{n})=1/\sqrt{n}$ with $\Pr(v=1)=1-1/\sqrt{n}$, for use in (5). This sampler is modeled on the ‘butcher-knife’ that Hesterberg, (1995) describes as an alternative to the jackknife. In some limited investigations it yielded very short intervals and it is not included in the numerical results of this paper.
3 Properties of the beta bootstrap $t$
Here we show that the beta bootstrap $t$ is second order accurate as $n\to\infty$. We also give conditions for $\mathbb{E}((t^{*})^{2}\!\mid\!x_{1},\dots,x_{n})<\infty$ that apply for small $n$. Finally, we compare the widths of $95\%$ bootstrap confidence intervals for the case with $n=2$ and $x_{1}\neq x_{2}$,
3.1 Second order accuracy
We show that the bootstrap $t$ with pseudo-counts $v_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}4\mathrm{Beta}(1/2,3/2)$ satisfies a critical condition from Hall and Mammen, (1994) for second order accuracy. They assume that $\mathbb{E}(x^{12})<\infty$ and do not impose the Cramér condition (1). Instead they show the moment conditions that must be satisfied. A Cramér condition is an additional requirement.
Their weights corresponding to $v_{i}$ are $W_{i}=nv_{i}/\sum_{i^{\prime}=1}^{n}v_{i^{\prime}}$. That is, $W_{i}=nw_{i}$. For second order accuracy, Hall and Mammen, (1994) state that it is necessary and sufficient to have both
$\mathbb{E}((W_{i}-1)^{2})=1+o(n^{-1/2})\quad\text{and}\quad\mathbb{E}((W_{i}-1)^{3})=1+o(1)$ (8)
along with $\mathbb{E}(W_{i})=1$ that they earlier assumed. We easily see that $\mathbb{E}(W_{i})=\mathbb{E}(\bar{W})$ by symmetry and $\bar{W}=1$ where $\bar{W}=(1/n)\sum_{i=1}^{n}W_{i}$. Therefore equation (8) gives the conditions on $v_{i}$ that remain to be established.
We consider IID pseudocounts $v_{i}>0$ with $\mathbb{E}(v)=1$, $\mathbb{E}(v^{2})=2$ and $\mathbb{E}(v^{3})=5$. Then $\mathbb{E}((v-1)^{2})=\mathbb{E}((v-1)^{3})=1$. Because $W_{i}$ are formed as ratios we have to account for their random denominator when verifying (8). For $\ell\geqslant 1$, we let $\mu_{\ell}=\mathbb{E}(v^{\ell})$.
Theorem 1.
Let $v_{i}$ be positive IID random variables with $\mu_{1}=1$, $\mu_{2}=2$ and $\mathbb{E}(v^{-\alpha})\leqslant M<\infty$ for some $\alpha>0$. If $\mu_{4+\epsilon}<\infty$ for some $\epsilon>0$, then
$\mathbb{E}(W_{i}^{2})$ $=2+\frac{10-2\mu_{3}}{n}+O(n^{-2}).$ (9)
If $\mu_{6+\epsilon}<\infty$ holds for some $\epsilon>0$ then
$\mathbb{E}(W_{i}^{3})$ $=\mu_{3}+\frac{9\mu_{3}-3\mu_{4}}{n}+O(n^{-3/2}).$ (10)
Proof.
See Appendix A.2. ∎
The beta bootstrap $t$ of Section 2.5 satisfies the conditions of Theorem 1. Because it has $\mu_{3}=5$, it then satisfies the conditions of Hall and Mammen, (1994).
3.2 Finiteness of some RMSLs
The beta bootstrap $t$ and the power bootstrap $t$ never give infinite length confidence intervals. The methods with $\Pr(w_{i}=0)>0$ give infinite length intervals for small $n$ even with continuously distributed data and they can have infinite length for substantially larger $n$ when the distribution of $x_{i}$ has atoms.
The finite lengths for the beta and power bootstrap $t$ intervals can still be large. Here we give conditions to ensure that $\mathbb{E}((t^{*})^{2}\!\mid\!x_{1},\dots,x_{n})<\infty$. That is a stronger condition than having a finite conditional quantile.
First we suppose that there are at least three distinct $x_{i}$ in the sample. This holds with probability one when $x_{i}$ have a continuous distribution. It also holds with high probability for some discrete distributions such as the $\mathrm{Pois}(1)$ distribution where all the bootstrap $t$ methods with $\Pr(v=0)>0$ gave an infinite RMSL for $n=19$ in Section 5.
Theorem 2.
Let $x_{1},\dots,x_{n}$ have $k\geqslant 3$ distinct observations, and let $t^{*}$ be the beta bootstrap $t$ statistic. Then $\mathbb{E}((t^{*})^{2}\!\mid\!x_{1},\dots,x_{n})<\infty$. The same holds for the power bootstrap $t$ statistic.
Proof.
See Appendix A.3
More generally if $v_{i}$ are IID $\mathrm{Beta}(a,(a^{2}+a)/(1-a))$, for $0<a<1$ and there are more than $1/a$ distinct observations $x_{i}$ in the sample, then $\mathbb{E}((t^{*})^{2}\!\mid\!x_{1},\dots,x_{n})<\infty$ holds by the argument behind Theorem 2.
Theorem 3.
Let there be two distinct observations among $x_{1},\dots,x_{n}$, each appearing at least three times. Then for the beta bootstrap $t$, $\mathbb{E}((t^{*})^{2}\!\mid\!x_{1},\dots,x_{n})<\infty$. The same holds for the power bootstrap $t$ statistic.
Proof.
See Appendix A.4. ∎
Similarly to the prior theorem, it suffices to have more than $1/a$ copies of each of the two distinct observations in order to get $\mathbb{E}((t^{*})^{2}\!\mid\!x_{1},\dots,x_{n})<\infty$.
3.3 Width when $n=2$
For $n=2$, with $x_{1}\neq x_{2}$, the proof of Theorem 3 leads to the expression
$t^{*}=\sqrt{2}\frac{w_{1}-1/2}{\sqrt{w_{1}(1-w_{1})}}\mathrm{sign}(x_{1}-x_{2}).$ (11)
Because $w_{1}=v_{1}/(v_{1}+v_{2})$, this allows us to compute the quantiles of $t^{*}$ and compare the different bootstrap $t$ methods to each other, to Student’s $t$ and to BCa. We can express all of the interval lengths as multiples of $|x_{1}-x_{2}|$.
For Student’s $t$ with $n=2$ we have $s^{2}=(x_{1}-x_{2})^{2}/2$. The usual Student’s $t$ interval reduces to
$\bar{x}\pm t^{0.975}_{(1)}\frac{s}{\sqrt{2}}=\bar{x}\pm t^{0.975}_{(1)}\frac{|x_{1}-x_{2}|}{2}.$
The $t_{(1)}$ distribution is the Cauchy distribution, and we get $t^{0.975}_{(1)}\doteq 12.71$. For the very benign case where $x\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(\mu,\sigma^{2})$, the exact confidence interval has width $12.71|x_{1}-x_{2}|$ when $n=2$.
For $n=2$, the weighted bootstrap $t$ interval widths are $\hat{\sigma}=|x_{1}-x_{2}|/2$ times the $0.95$ quantile of the distribution of $2|w_{1}-1/2|/\sqrt{w_{1}(1-w_{1})}$. That is, the width is the $0.95$ quantile of $|w_{1}-1/2|/\sqrt{w_{1}(1-w_{1})}$. For the usual multinomial bootstrap $t$, $w_{1}$ takes values 0, $1/2$ and $1$ with probabilities $1/4$, $1/2$ and $1/4$ respectively. Then the width takes values 0 and $\infty$ with probability $1/2$ each and the $95$’th percentile of this distribution is $\infty$. The Poisson bootstrap $t$, double-or-nothing bootstrap $t$ and half-sampling bootstrap $t$ all give infinite intervals for $n=2$ with $x_{1}\neq x_{2}$.
For the BCa with $n=2$, the formula in Section 2.3 yields $x_{(1)},x_{(1)}]$ where $x_{(1)}=\min(x_{1},x_{2})$. While $\Pr(\bar{x}^{*}=\bar{x})$ is negligible for modestly large $n$, it equals $1/2$ for $n=2$. If we modify equation (4) to put half weight on cases where $\bar{x}^{*}=\bar{x}$, then we get a 95% interval of $x_{(1)},x_{(2)}]$ which is a truer reflection of how BCa operates. The numerical results in this paper use that modification. When $\bar{x}^{*}=\bar{x}$ arises because the same values were averaged in both cases, they will also be equal in their floating point representations.
Table 2 shows width factors for various ACIs when $n=2$. The bootstrap ones were computed numerically by sampling values of $w_{1}=v_{1}/(v_{1}+v_{2})$. What we see there is that most of the bootstrap ACIs we consider are either narrower than Student’s $t$ intervals or have infinite length when $n=2$. It is reasonable for a nonparametric interval to be wider than one based on a Gaussian data assumption. Of the methods in Table 2, only the power bootstrap $t$ has intervals wider than Student’s $t$ but of finite length.
[…]
Multinomial $\infty$
Power 18.60
Student 12.71
Beta 10.78
Bayesian 03.04
Log normal 01.43
Mammen 01.12
BCa  111
Table 2: This table shows widths of $t$-based confidence intervals at nominal level 95%, when $n=2$ as a multiple of $|x_{1}-x_{2}|$. Multinomial is the usual bootstrap $t$. Student is the usual Student’s $t$ interval. The others are weighted bootstraps.
4 Towards a common task framework for ACIs
The considerable progress in black box prediction and natural language processing has been attributed to the common task framework epitomized by DARPA’s challenge problems. That framework includes standard benchmark problems on which to compare objectively measurable performance such as predictive accuracy. In those settings, state of the art performance need not always be accompanied by a full explanation of why a method works well. See Donoho, (2024) who points to a lecture by Mark Liberman for the term.
There is no widely accepted measure of performance for ACIs, so there is no accepted benchmark with which to judge state of the art performance. The main difficulty is that we have two competing objectives: attaining something close to the nominal coverage level, and having intervals that are as short as possible. Sometimes one method has better coverage than another along with shorter intervals. Then that other method is inadmissible. Such domination is rare among seriously considered methods, and they span a large Pareto frontier.
Confidence intervals are similar to prediction intervals and also to quantile elicitation, where there are established criteria. Binary classification has an established $F_{1}$ score to make a similar tradeoff between precision and recall. There are also decision theoretic methods to measure performance of confidence intervals. We see below how none of these concepts provide a good benchmark for ACIs.
4.1 Calibration
For both confidence and prediction intervals, we provide an ideally short interval $[L,U]$ and there is a desired nominal coverage probability $\alpha$. Here $L=L(X)$ and $U=U(X)$ where $X$ represents all the data. In a prediction interval we seek $\Pr(L\leqslant Y\leqslant U\!\mid\!X)=\alpha$ for random $Y$ conditionally on $X$. In a confidence interval we seek $\Pr(L(X)\leqslant\theta\leqslant U(X))$ for a non-random quantity $\theta$ that is a function of the distribution of $X$, such as the mean $\mu$. The probability is under randomess in $X$.
Prediction intervals can be calibrated on an exchangeable set of problem instances in order to get close to the nominal coverage. For a recent survey of calibration based on conformal inference, see Angelopoulos and Bates, (2023). There, in an exchangeable sequence of $N$ prediction problems it is possible to calibrate so that the coverage of the $n$’th one is between $1-\alpha$ and $1-\alpha+1/(n+1)$. We can compare well-calibrated prediction intervals by preferring shorter ones using mean length or some other aspect of the distribution of their interval lengths.
Confidence intervals are harder to calibrate than prediction intervals. We don’t generally observe the true parameter value. We can calibrate them numerically for specific hypothesized distributions of $x_{i}$. We usually lack a way to show that our problem is exchangeable within some set of hypothesized problems that we can calibrate.
4.2 Elicitation and the interval score
Continuing the comparison between confidence intervals and prediction intervals, the one-sided version of a prediction interval is a quantile $Q^{\alpha}$ satisfying $\Pr(x\leqslant Q^{\alpha})=\alpha$, which for simplicity we will assume is unique. Quantiles minimize the expected value of a piecewise linear ‘pinball’ score function, familiar from quantile regression (Koenker,, 2005), given by
$S_{\alpha}(\theta;x)=(x-\theta)\bigl{(}\alpha-1_{x<\theta}\bigr{)}.$ (12)
The minimizer over $\theta$ of $\mathbb{E}(S_{\alpha}(\theta;x))$ is the $\alpha$ quantile of $x$ very generally. Gneiting and Raftery, (2007) describe a requirement that either $\mathbb{E}(\max(x,0))<\infty$ or $\mathbb{E}(\max(-x,0))<\infty$, both of which are satisfied if $\mathbb{E}(|x|)<\infty$.
The literature on scoring mixes settings where the score function is a utility to be maximized with those where it is a loss to be minimized. We will use losses, translating some of the results for utilities to this case.
A scoring function $S$ is consistent for a real-valued functional $T$ of the distribution $F$ relative to the class $\mathcal{F}$ of distributions if $\mathbb{E}_{F}(S(\theta,x))\geqslant\mathbb{E}_{F}(S(T(F),x))$ for all $F\in\mathcal{F}$ and all $\theta\in\mathbb{R}$ when $x\sim F$. It is strictly consistent if equality holds only when $\theta=T(F)$. If $T(\cdot)$ has a strictly consistent scoring function, then it is elicitable (relative to the class $\mathcal{F}$). For the distinction between consistent score functions and the closely related proper score functions, see Gneiting, (2011).
Quantiles are elicitable because $S_{\alpha}$ above is a consistent score function. This function can be used to elicit an expert’s opinion of the quantile of $x$. When $\mathbb{E}(\cdot)$ is with respect to that expert’s opinion of the distribution of $x$, then their best choice is to answer with their true opinion about $Q^{\alpha}$ instead of, for example, cautiously over-estimating the probability of some undesired outcome (such as rain or financial loss or illness). Any such attempts to game the problem can only worsen the expert’s score.
We can form a prediction interval from two quantiles such as $Q^{\alpha_{j}}$ for $j=1,2$. For a two-sided symmetric prediction interval at level $\alpha$, we can take $\alpha_{1}=(1-\alpha)/2$ and $\alpha_{2}=1-\alpha_{1}$ and sum the scores getting
$S_{\alpha_{1}}(x,\theta_{1})+S_{\alpha_{2}}(x,\theta_{2})$ $=(\alpha_{1}-1_{x\leqslant\theta_{1}})(x-\theta_{1})+(1-\alpha_{1}-1_{x\leqslant\theta_{1}})(x-\theta_{2})$
$=\alpha_{1}(\theta_{2}-\theta_{1})+1_{x\leqslant\theta_{1}}(\theta_{1}-x)+1_{x>\theta_{2}}(x-\theta_{2}).$
The interval length between the upper and lower quantiles is $\theta_{2}-\theta_{1}$. If we divide the score above by $\alpha_{1}$, we get the interval score
$\theta_{2}-\theta_{1}+\frac{1_{x<\theta_{1}}}{\alpha_{1}}(\theta_{1}-x)+\frac{1_{x>\theta_{2}}}{\alpha_{1}}(x-\theta_{2})$ (13)
due to Winkler, (1972). The above derivation uses $1_{x\leqslant\theta_{1}}(\theta_{1}-x)=1_{x<\theta_{1}}(\theta_{1}-x)$. Equation (13) is equation (43) of Gneiting and Raftery, (2007). Their $\alpha$ is our $1-\alpha$, so for the default coverage setting, our $1/\alpha_{1}=40$ is their $2/\alpha=40$.
When scoring confidence intervals $[L,U]$ for the mean $\mu$, if we reverse the notion of what is random, as if $\mu$ were random with estimated quantiles $L$ and $U$, then the Winkler score (13) yields
$U-L+\frac{1_{\mu<L}}{\alpha_{1}}(L-\mu)+\frac{1_{\mu>U}}{\alpha_{1}}(\mu-U).$ (14)
This measure combines the length ULU-L of the interval with some non-coverage penalties that grow with the distance between $\mu$ and the interval $[L,U]$. This measure was proposed in the masters thesis Hofer, (2022) and used in Hofer and Held, (2022) to score confidence intervals for the mean $\pi$ of a Bernoulli distribution. They find that Wald intervals perform badly, and for $\pi$ not near 0 or $1$, the best results were from a Wilson confidence interval or a Bayesian credible interval using a uniform prior.
Scoring a confidence interval the way we would score a prediction interoval involves reversing the role of what is fixed with what is random. Statisticians are very used to this when alternating between frequentist and Bayesian viewpoints of confidence/credible intervals. So it is surprising that the proposal in Hofer and Held, (2022) is so recent.
An upper confidence interval function would be a statistic $T_{n}(x_{1},\dots,x_{n})$ with the property that
$\Pr\bigl{(}\mathbb{E}(x_{1})\leqslant T_{n}(x_{1},\dots,x_{n})\bigr{)}=\alpha$
for $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$. By the result of Bahadur and Savage, (1956) we know that no non-trivial such $T_{n}$ exists. Perhaps we should not expect to elicit a statistical quantity that does not exist. Some quantities that are closely related to confidence intervals are known to not be elicitable. Frongillo and Kash, (2014) call the interval $(a,b)$ a 90% confidence interval for $x$ if $\int_{a}^{b}p(x)\,\mathrm{d}x=0.9$. That does not correspond to our usage of the term ‘confidence interval’. It is a prediction interval. They go on to show that the entire set of such pairs $(a,b)$ is not elicitable, nor is the pair with the smallest width $\omega=b-a$, nor is $\omega$ itself elicitable. For more in this direction, see Brehmer and Gneiting, (2021). From here on, we assume that there is no consistent score function for non-trivial confidence intervals.
4.3 Decision theory
There has been some work on decision theoretic frameworks for confidence intervals and set estimation more generally. Good starting points for that literature are Casella and Hwang, (1991) and Casella et al., (1994). Those works are for a parametric setting, which is understandable due to the non-existence of fully nonparametric confidence intervals for the mean. The most prominent loss functions take a form equivalent to
$\lambda\times(U-L)-1\{L\leqslant\mu\leqslant U\}$ (15)
for some $\lambda>0$. Authors vary in whether the factor $\lambda$ is applied to the interval length or the coverage indicator. Then the risk is
$\lambda\times\mathbb{E}(U-L)-\Pr(\mu\not\in[L,U]),$ (16)
trading off expected length and coverage probability. The great difficulty is in choosing the value for $\lambda$. If we are comparing methods that achieve $\Pr(\mu\not\in[L,U])=\alpha$ then any $\lambda>0$ leads us to prefer shorter mean length. When we have to compare two or more imperfectly calibrated methods, as we must for ACIs, then it is very hard to choose the constant $\lambda$.
A more serious difficulty is that if we fix $\lambda$ and then minimize risk, we are also effectively choosing the level $\alpha$ for our interval. When we are reporting a confidence interval or plotting error bars we want to directly specify and communicate the nominal level. It would not be satisfactory for the coverage level $\alpha$ corresponding to some $\lambda$ to be different from our nominal $\alpha$. It would be even worse to communicate intervals whose nominal level was unknown.
Casella et al., (1994) describe a paradox told to them by James Berger. When using the loss (15) on Gaussian data, the usual Student’s $t$ interval can be dominated by a strategy that truncates the interval to length 0 when the sample standard deviation $s$ is large enough. The paradox can be removed through the use of an increasing scale function $S:[0,\infty)\to[0,1]$ and then using
$S(U-L)-1\{L\leqslant\mu\leqslant U\},$ (17)
where $S$ could incorporate a scaling factor like $\lambda$.
For ACIs, the coverage error is typically $O(1/n)$ while the mean length is typically $O(1/\sqrt{n})$. Then if $\lambda$ is reasonable for one value of $n$ it is not suitable for other values of $n$. We can make the choice of $\lambda$ less dependent on $n$ by either penalizing the mean squared length, or by using
$\lambda\frac{U-L}{\sqrt{n}\sigma}-1\{L\leqslant\mu\leqslant U\}.$ (18)
Scaling by $\sigma$ makes the performance invariant to rescaling the data $x_{i}$ to $cx_{i}$ for $c>0$ when, as is usual, coverage is also invariant to such rescaling. In numerical examples, we know $\sigma$ and can then use it in comparisons of methods.
It is not obvious that an interval becomes better when its coverage level exceeds the nominal level. Over-estimating the coverage of the interval $[L,U]$ necessarily under-estimates the coverage of $(-\infty,L)\cup(U,\infty)$. For 95% coverage, criteria like
$\lambda\frac{\mathbb{E}((U-L)^{2})^{1/2}}{\sigma\sqrt{n}}-|\Pr(\mu\not\in[L,U])-0.95|,$ (19)
or
$\lambda\frac{\mathbb{E}((U-L)^{2})^{1/2}}{\sigma\sqrt{n}}-\min\bigl{(}0.95,\Pr(\mu\not\in[L,U])\bigr{)}$ (20)
explicitly account for the target coverage level, unlike (16). Using the root mean squared length (RMSL) in these criteria penalizes methods that produce highly variable interval lengths.
The measure in (19) penalizes both under-coverage and over-coverage. The measure in (20) does not explicitly penalize over-coverage, but does not also reward it. Confidence interval constructions usually have a parameter such as $t_{(n-1)}^{\alpha}$ in Student’s $t$ that can be raised or lowered to trade off length and coverage. Methods that over-cover are then methods that might otherwise have had nominal coverage at lower length, e.g., smaller $t^{\alpha}$. Their over-coverage is then already penalized by that extra interval length. Neither of the above criteria can be written as values of expected loss functions. In that sense they are only ‘quasi-risks’, not risks. In any of the criteria (16), (19) and (20) it remains unclear how one should choose $\lambda$.
4.4 The $F_{1}$ score
In binary classification problems, some methods may have better precision while others have better recall. These two desiderata conflict in a very similar way to how the criteria for ACIs do. The $F_{1}$ score is commonly used to combine them into one criterion. The context is that of predicting a binary label $y\in\{0,1\}$ where items with $y=1$ are rare and interesting to discover, while $y=0$ is the norm. This is similar to the problem of choosing a confidence interval where we must trade off two competing criteria. And yet, the $F$-score described next is a very popular, though not universally popular, way to do it. An early context was information retrieval (Van Rijsbergen,, 1979) where $y=1$ describes a relevant document and earlier uses were in biology (Dice,, 1945; Sørensen,, 1948).
[…]
$F_{1}=\frac{1}{(1/p+1/r)/2}=\frac{2pr}{p+r}$
and it is widely used as a single measure of classifier accuracy. For some debates about its suitability see Hand et al., (2021) and for a listing of many alternatives, see Powers, (2020). The $F_{1}$ score is preferred to the overall accuracy $\Pr(\hat{y}=y)$ because the latter can be very high for the completely uninformative rule with $\Pr(\hat{y}=0)=1$. The more general $F_{\beta}$ score given by
$F_{\beta}=\frac{\beta^{2}+1}{\beta^{2}/r+1/p}$
puts more emphasis on recall when $\beta>1$ and more emphasis on precision when $\beta<1$.
We can draw an analogy between binary classification and confidence intervals. The true value $\mu$ is the one real number that we wish to recover. That makes coverage probability the analogue of recall. Taking $\tilde{r}=\max(\Pr(L\leqslant\mu\leqslant U),0.95)$ allows us to incorporate a specific nominal coverage level, which we were not able to do with the Winkler score. This criterion is designed to reward absolute coverage (up to a limit) instead of penalizing coverage error.
For precision, a shorter interval corresponds to greater precision. We want something proportional to $1/\mathrm{RMSL}$. We choose $\tilde{p}=\sigma/(\sqrt{n}\,\mathrm{RMSL})$. The scaling by $\sigma$ makes the criterion invariant to rescaling $x_{i}$ and this is appropriate because $\tilde{r}$ is invariant. Where coverage error decreased as $1/n$, absolute coverage approaches the nominal level and our scaling by $\sqrt{n}$ makes $\tilde{p}=O_{p}(1)$. With these choices
$\tilde{F}_{\beta}=\frac{\beta^{2}+1}{\beta^{2}/\max(\Pr(L\leqslant\mu\leqslant U),0.95)+\sqrt{n}\mathrm{RMSL}/\sigma}.$ (21)
To get a $\tilde{F}_{\beta}$ score suitable for ACIs we would need a consensus choice for $\beta$.
4.5 Some other criteria
L’Ecuyer et al., (2023) prioritized coverage over length. For a nominal 95% coverage any method delivering below 94% coverage would be considered a failure. If a method did not cover the true mean at least 927 times out of 1000, then it was deemed to have failed. A false positive would happen with less than 4% probability, by the binomial distribution. That large study of randomized quasi-Monte Carlo integration featured 2400 tasks from 6 integrands, 5 RQMC methods, 4 integral dimensions, 5 RQMC sample sizes and 4 choices for the number of RQMC replicates to use. The percentile bootstrap failed 1698 times, the bootstrap $t$ failed 81 times, and the plain Student’s $t$ interval failed only 3 times.
The RQMC setting is special. Recent literature is showing that the integral estimates are nearly symetrically distributed with large kurtosis. For one kind of RQMC, the diverging kurtosis follows from results in Pan and Owen, (2023) while the modest skewness has been studied in Pan and Owen, (2025). It is well known that the standard Student ACIs attain high coverage, often over the nominal level, for data that come from a symmetric distribution with very heavy tails. See Miller, (1986), Efron, (1969) and Cressie, (1980).
Owen, (1988) compared nonparametric ACIs for the mean. Portions of that work appeared in Owen, (1992). In one of the comparisons each method chose an ACI and then each other method selected its ACI using the same length that the first method chose. That allows us to compare methods by a measure of length adjusted coverage. In that framework, the empirical likelihood intervals generally did best. They were not however as well calibrated as bootstrap $t$. Improving the calibration of empirical likelihood intervals is outside the scope of this article.
Wolfowitz, (1950) described an interesting way to modify the length of a confidence interval in a loss function. When an interval does not cover $\mu$, he extends it until it does and then uses the length of the resulting interval $\min(U,\mu),\max(L,\mu)]$. In some numerical work this did not make much difference. That can be explained by non-coverage being relatively rare and the distance between the interval and $\mu$ being $O_{p}(n^{-1/2})$.
4.6 Illustration
To illustrate the difficulties in quantifying ACI quality, consider ACIs for the mean of the $\mathrm{Exp}(1)$ distribution based on $x_{1},\dots,x_{10}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Exp}(1)$. This distribution is only a moderately difficult one from Section 5. The sample size $n=10$ is near the middle of the range $2\leqslant n\leqslant 20$ considered there.
The true mean is $\mu=1$ and the mean coverage of any given ACI is now a $10$ dimensional integral. We can approximate that integral using 10,00010{,}000 vectors $X\in\mathbb{R}^{10}$ with IID $\mathrm{Exp}(1)$ components. The results are in Table 3 for bootstrap methods based on $B=2000$ resamples, sorted in decreasing order by coverage. If we go by coverage and average length, then all of the methods are on the Pareto frontier, except for the half sampling bootstrap $t$, which has greater mean length and lower coverage than the Poisson bootstrap $t$.
Method Cover Length Length2 Win95 Win99 $\tilde{F}_{1}$
[…]
Poisson 0.9612 2.26 8.07 2.49 3.40 0.200
Half Sample 0.9577 2.37 9.37 2.62 3.60 0.187
Multinomial 0.9415 1.87 5.40 2.22 3.62 0.238
Power 0.9400 1.62 3.32 1.98 3.40 0.293
Beta($1/2,3/2$) 0.9311 1.52 2.90 1.94 3.59 0.311
Student’s t 0.9002 1.33 2.07 1.96 4.50 0.357
Exponential 0.8935 1.23 1.84 1.90 4.58 0.374
BCa 0.8772 1.17 1.64 1.92 4.95 0.392
Mammen 0.8681 1.12 1.43 1.96 5.36 0.413
Lognormal 0.8405 1.01 1.20 2.04 6.17 0.443
Table 3: This table describes 9595% ACIs for a sample of $n=10$ draws from $\mathrm{Exp}(1)$. All the methods are weighted bootstrap $t$ except for BCa and Student’s $t$. The columns are mean coverage, mean interval length, mean squared interval length, two Winkler scores, one at level $0.95$ and one at level $0.99$, and $\tilde{F}_{1}$ from (21).
If we use the Winkler score for level $0.95$ then the best score comes from a bootstrap $t$ using Rubin’s Bayesian bootstrap, i.e., exponential weights. It does not seem reasonable that a method with 89.3589.35% coverage should be best when other methods attain 93% to 96% coverage. Those other methods use longer intervals, but they do not appear to have drastically greater length.
Sometimes experts making a prediction might make a prediction that is more cautious than their true belief if the score function is not proper. In a setting with no proper score function like ACIs appear to be, we might judge that the 95% Winkler score is too lenient on coverage (versus length) and compensate by scoring nominally 95% ACIs by a higher coverage goal such as 99%. By that measure we would favor the bootstrap $t$ with power weights. It raises the coverage from just over 89% to 94% while increasing the mean length by a factor of $1.62/1.23\doteq 1.32$ and the mean squared length by $3.32/1.84\doteq 1.80$.
Poisson weights attain essentially the same score as power weights do here. They do so by getting credit for coverage that goes beyond even 95%. While it may or may not be reasonable to penalize a method for covering more than the nominal level, it does not seem reasonable to credit it for covering more than the nominal level. This breaks the Winkler score tie in favor of the power bootstrap $t$.
If we insist on at least 95% coverage and then minimize mean or mean squared length subject to that constraint, then Poisson weights did best for this case. If we go by the largest values of $\tilde{F}_{1}$ then the favored methods would be the ones that most severely under-cover $\mu$. In other words, we are not fortunate enough to have a setting where $\beta=1$ provides a reasonable default. For other values of $\beta$ $\tilde{F}_{\beta}$ generally favors either the longest or the shortest measures. It is essentially picking one criterion and neglecting the other.
The author’s subjective judgement of the results in Table 3 is as follows. The multinomial, power and beta bootstrap $t$ methods are all reasonable choices attaining nearly nominal coverage without excessive length. The bottom five methods, from Student’s $t$ to the lognormal bootstrap $t$ have coverage levels that are unacceptably low by comparison. While the top three methods cover even more than the target level, they do so with unreasonable extra length.
This case also illustrates how some bootstrap $t$ methods, such as those using lognormal or Mammen weights can cover even less often than BCa does. At the same time, three of the bootstrap $t$ methods cover even more than the usual multinomial bootstrap $t$ does. Those are the ones that have even greater probabilities of assigning a weight of zero to an observation.
5 Test cases
We can examine some specific distributions to find the resulting coverage probabilities and interval lengths for different bootstrap methods. We begin with the asymptotic theory from Hall, (1988) because that lets us select distributions to provide quite different tests of the ACIs. Hall’s asymptotic theory is based on moments with assumed finite eighth moments and a Cramér condition. So we include some distributions with unbounded fourth moment and also a discrete distribution for which the Cramér condition does not hold.
5.1 Asymptotic coverage
Table 1 of Hall, (1988) provides some asymptotic coverage formulas for bootstrap ACIs for the mean of $n$ IID observations as the sample size $n\to\infty$. Owen, (2025) added a formula for Student’s $t$ method. The method of Hall, (1988) for Normal theory used $\hat{\sigma}$ and quantiles of $\mathcal{N}(0,1)$ instead of $s$ and $t_{(n-1)}$ quantiles.
The coverage errors for 95% confidence intervals are $\Pr(L\leqslant\mu\leqslant U)-0.95$, so negative values indicate undercoverage. The asymptotic formulas involve the skewness $\gamma=\mathbb{E}((x-\mu)^{3})/\sigma^{3}$ and the kurtosis $\kappa=\mathbb{E}((x-\mu)^{4})/\sigma^{4}-3$. The coverage errors for two-sided 95% confidence intervals are
Normal theory: $\quad(1/n)\varphi(z^{0.975})\bigl{[}\phantom{-}0.14\kappa-2.12\gamma^{2}-3.35\bigr{]}+O(1/{n^{2}}),$
Student’s $t$: $\quad(1/n)\varphi(z^{0.975})\bigl{[}\phantom{-}0.14\kappa-2.12\gamma^{2}\,\,\,\phantom{-3.35}\bigr{]}+O(1/{n^{2}}),$
Percentile: $\quad(1/n)\varphi(z^{0.975})\bigl{[}-0.72\kappa-0.37\gamma^{2}-3.35\bigr{]}+O(1/{n^{2}}),$
Bootstrap $t$: $\quad(1/n)\varphi(z^{0.975})\bigl{[}-2.84\kappa+4.25\gamma^{2}\,\,\,\phantom{-3.35}\bigr{]}+O(1/{n^{2}}),\quad\text{and}$
BCa: $\quad(1/n)\varphi(z^{0.975})\bigl{[}-2.63\kappa+3.11\gamma^{2}-3.35\bigr{]}+O(1/{n^{2}}),$
where $\varphi$ is the $\mathcal{N}(0,1)$ probability density function.
The error expressions above help to explain why the bootstrap $t$ is often so effective. It has the largest positive coefficient for $\gamma^{2}$ of all the methods in Table 1 of Hall, (1988) and it is the only bootstrap method with a zero intercept instead of 3.35-3.35.
Plugging $n=18$ into Hall’s formula gives very close the coverage error averaged over $16\leqslant n\leqslant 20$ in Table 6 of Owen, (1988) for all distributions but one. That one was the lognormal distribution where the formula gave a coverage error so low that it corresponded to negative coverage for the ACI.
Some of the gaps in Hall’s analysis can be filled in by parts C, D and E of Theorem 1 of Singh, (1981) which assume that $\mathbb{E}(|x-\mu|^{3})<\infty$. Then coverage errors are $o(n^{-1/2})$ in the non-lattice case and $O(n^{-1/2})$ in the lattice case.
5.2 Computations
Each sampling distribution was represented in R=10, [^0] R=10{,}000 data sets of sizes $2\leqslant n\leqslant 20$. All of the ACI methods were given the same 10,00010{,}000 samples of each size so they all received the same problem difficulty. Furthermore the samples of size $n+1$ always contained the points from the samples of size $n$. That technique gives more precise estimates of how length or coverage vary with $n$. For R=10, [^0] R=10{,}000 replicates, and a method with coverage probability $0.95$, the observed number of of coverages has $0.005$ quantile 9443 and $0.995$ quantile 9555. Then the coverage proportion has about 99.999.9% probability of being within $0.57$ percentage points of nominal. For a method with coverage 8080% comparable intervals are roughly $0.80\pm 0.01$. Every bootstrap was based on $B=2000$ resamplings.
Since there is no measure to elicit overall ACI quality and no generally accepted way to trade off length versus coverage, we plot a measure of interval length versus coverage in order to visualize the Pareto frontier. Based on the discussion in Section 4.3, interval length is measured by the root mean squared length (RMSL) divided by $\sqrt{n}$ so that it approaches zero as $O(1/n)$ which is the same rate that two-sided coverage typically approaches the nominal level. We also divide the RMSL by $\sigma=\mathrm{var}(x)^{1/2}$ to make the results for different distributions more nearly comparable.
[…]
The methods can be clustered into a few groups. The bootstrap $t$ methods with the greatest values of $\Pr(v_{i}=0)$ are those with double-or-nothing, half sampling and Poisson weights. Those methods generally have the longest intervals and the highest coverage. They also have the most frequent occurences of infinite length intervals. They are plotted in red/brown/orange colors, with the more reddish color for the longer intervals. The second group is made up of the usual multinomial bootstrap $t$ as well as the ones with power and beta distributed weights have lengths and coverage levels lower than the first group. They are colored by shades of green, with brighter green for the methods with generally longer intervals. A third group with even shorter intervals and generally very low coverage is made up of the BCa along with bootstrap $t$ with Mammen weights and exponential (i.e., Bayesian bootstrap) weights. They are plotted in shades of blue and the ordering among them varies with the setting. ACIs based on Student’s $t$ distribution do not fit cleanly into any of those three groups. They appear in different groups depending on the distribution of $x$, and are shown in black. The bootstrap $t$ using lognormal weights have extremely short intervals and low coverage. They are included only to show that it is possible to have weighted bootstrap $t$ intervals with lower coverage (and length) than BCa intervals have. They are shown in gray.
5.3 Benign/Gaussian data
A very benign setting for nonparametric ACIs arises when the $x_{i}$ have a unimodal density function with at most mild skewness and light tails. As an archetypal example of such a distribution, we consider $x_{i}\sim\mathcal{N}(0,1)$. Of course, in practice a user would not know that the data were from such a parametric model. We want to see how methods do on this case because a method that does poorly in even such a simple setting is not a good choice overall.
[…]
Figure 1 shows estimates of RMSL and coverage for Gaussian data. There are two views, one for $10\leqslant n\leqslant 20$ and one for $2\leqslant n\leqslant 10$. The methods cluster according to the earlier described groupings. The double-or-nothing bootstrap has by far the longest intervals and it covers far more than the nominal amount. Mammen, exponential and BCa intervals are all very short and under-cover significantly, even on this easy case. The multinomial, power and beta bootstrap $t$ intervals are intermediate with approximately nominal coverage and are close to the Student $t$ method which has exact coverage for this problem.
The bottom of the curves in the left panel shows that for $n=20$ there is a near perfect monotone relationship between the coverage level and the RMSL. The one exact method, Student’s $t$ interval, has a very slightly better RMSL than the multinomal and Beta bootstrap $t$ methods with nearby coverage levels.
From the right panel we can see that the double-or-nothing bootstrap only had a finite RMSL for $n\geqslant 9$. This simply means that at least one of the 10,00010{,}000 ACIs it produced had infinite length. An analysis in Appendix C shows that for distinct $x_{i}$ and $B\to\infty$ that a finite length ACI will happen for $n\geqslant 9$ but not necessarily for $n\leqslant 8$. The Poisson ACIs were only finite for $n\geqslant 7$ which also matches the cutoff in Appendix C. When $n$ is just barely large enough to assure a finite ACI length, then the RMSL can be very heavy-tailed. The plot does not show every finite RMSL in order to better show the bulk of them.
$n=2$ $n=3$ $n=4$
Method Cover RMSL Cover RMSL Cover RMSL
Power $0.9650$ $26.55$ $0.9563$ $7.66$ $0.9542$ $4.09$
Student’s $t$ $0.9493$ $17.94$ $0.9501$ $4.98$ $0.9517$ $3.17$
Beta $0.9389$ $15.31$ $0.9356$ $5.76$ $0.9398$ $3.48$
Table 4: Coverage and unscaled RMSL for small sample sizes on Gaussian data.
The power and beta bootstrap $t$ methods do very well for the Gaussian data even with tiny $n$. Table 4 shows the coverage and RMSLs for $2\leqslant n\leqslant 4$ for the power and beta bootstrap $t$ methods along with the value for Student’s $t$. The coverage errors are much closer to nominal there than we see for much larger $n$ on more difficult distributions.
5.4 Skewed/exponential data
Skewness is one of the main reasons to switch from Student’s $t$ distribution to some kind of bootstrap distribution. Our archetypal example for mildly skewed data has $x_{i}\sim\mathrm{Exp}(1)$. It has skewness $2$. There is no universal rule for what amount of skewness is mild and what is severe. Some authors consider skewness in [1,1][-1,1] to be mild and skewness $2$ to already be severe. Others consider skewness of $2$ to not be extreme. Where Hesterberg, (2015) recommends $n\geqslant 4815$ for Student’s $t$ test he finds that the bootstrap $t$ with $n\geqslant 101$ meets his stringent conditions.
Figure 2: Same as Figure 1 except that $x_{i}\sim\mathrm{Exp}(1)$.
Figure 2 shows results for $x\sim\mathrm{Exp}(1)$. Now Student’s $t$ clusters with BCa and the Mammen and exponential bootstraps, instead of with the beta, power and multinomial bootstrap $t$. Student’s $t$ and BCa  under-cover significantly. The multinomial, power and beta bootstraps slightly under-cover. The power version has comparable coverage to the multinomial over $10\leqslant n\leqslant 20$ with somewhat smaller RMSL.
Half sampling, Poisson and double-or-nothing bootstraps over cover and have much greater length. The double-or-nothing bootstrap has infinite RMSL for $n\leqslant 8$.
[…]
As noted above the standard Student ACIs attain high coverage on symmetric heavy tailed data. Our archetypal example for this case is $x_{i}\sim t_{(4)}$. This distribution is symmetric with skewness zero and infinite kurtosis. As a result the asymptotic coverage results from Hall, (1988) do not apply for it, but Theorem 1C of Singh, (1981) shows that coverage errors are $o(n^{-1/2})$ for the bootstrap $t$ because $\mathbb{E}(|x|^{3})<\infty$ and the distribution is non-lattice.
Figure 3: Same as Figure 1 except that $x_{i}\sim t_{(4)}$.
Student’s $t$ ACIs are the best by far here in line with well known theory. All of the bootstrap $t$ method under-cover except for the double-or-nothing version which has infinite RMSL for $n\leqslant 8$. BCa barely attains over 90% coverage by $n=20$.
[…]
Here we consider light-tailed symmetric data, using $\mathbb{U}(0,1)$ as the example. Figure 4 shows the results. Student’s $t$ does best which is not surprising for symmetric data. For $n\geqslant 10$, most methods significantly over-cover the true mean. The Bayesian bootstrap (exponential weights) and BCa under-cover but come quite close to the nominal level.
Figure 4: Same as Figure 1 except that $x_{i}\sim\mathbb{U}(0,1)$.
5.7 Severe/lognormal setting
A severe setting has a very large kurtosis but is also not symmetric enough for Student’s $t$ ACIs to do well. For this case, we consider the lognormal distribution $x_{i}\sim\exp(\mathcal{N}(0,1))$. None of the bootstrap methods in Owen, (1988) were able to get much more than 90% coverage by $n=20$ for this distribution when the nominal rate was 95%. The predicted coverage using the formulas of Hall, (1988) in Section 5.1 was actually negative because of the extremely large kurtosis. This distribution has
$\kappa$ $=\exp(4)+2\exp(3)+3\exp(2)-6\doteq 110.93\quad\text{and}$
$\gamma$ $=(\exp(1)+2)\sqrt{\exp(1)-1}\doteq 6.18.$
Figure 5: Same as Figure 1 except that $x_{i}\sim\exp(\mathcal{N}(0,1))$.
This lognormal data is very heavily skewed and has an enormous kurtosis. For $10\leqslant n\leqslant 20$, the double-or-nothing bootstrap has somewhat over nominal coverage. It is extremely hard for any of the other methods to cover well. Student’s $t$ ACIs severely undercover and are clustered with BCa and the Bayesian bootstrap. The best coverage levels only arose for methods with $\Pr(v_{i}=0)>0$. For the larger sample sizes in this window the power bootstrap $t$ had shorter intervals and nearly the same coverage as the multinomial bootstrap $t$.
For the very small $n\in[2,10]$, the shortest methods barely make it into the plotting region with a coverage cutoff near $0.8$. We also see some non-monotonicity in the coverage levels.
5.8 Discrete mildly skewed data: $\mathrm{Pois}(1)$
Data with a discrete distribution can generate numerous tied values among the $x_{i}$. The presence of such ties makes $\hat{\sigma}^{*}=0$ much more common. This is especially problematic for methods that also allow $v_{i}=0$.
Another difficulty arises when the distribution of $x_{i}$ only takes integer values, or more generally, only takes values within some arithmetic sequence in $\mathbb{R}$. In this case the distribution $F$ does not satisfy a Cramér condition that was a key assumption in the asymptotic analysis of Hall, (1988) and others. There is thus much less known about coverage properties even for large $n$. For count data, with small $n$ and integer counts $w_{i}$, it can easily happen that $\bar{x}^{*}=\bar{x}$ and then the distribution of $t^{*}$ has an atom at 0.
Our example for this setting is $x_{i}\sim\mathrm{Pois}(1)$. It has mild skewness and kurtosis both equal to $1$. Second order accuracy does not hold for such data.
Figure 6: Same as Figure 1 except that $x_{i}\sim\mathrm{Pois}(1)$.
Results for Poisson data are shown in Figure 6. When the top (or bottom) $2.5\%$ of $t^{*}$ values are infinite the resulting ACI has infinite length. All of the double-or-nothing and half sampling ACIs had infinite RMSL for $2\leqslant n\leqslant 20$. For the Poisson bootstrap and the multinomial bootstrap the first finite RMSL was for $n=20$. The clear lesson is that for discrete data, we should avoid bootstraps with $\Pr(v_{i}=0)>0$.
For $10\leqslant n\leqslant 20$, the beta bootstrap slightly over covered. Among the under-covering methods, the Mammen bootstrap $t$ did the best. For $n\leqslant 10$ the attained coverage levels of several methods are not monotone in $n$ although the RMSLs are monotone.
5.9 Discrete skewed data/geometric data
We also consider a geometric distribution equal to the number of failures seen prior to the first success in a sequence of Bernoulli trials with success probability $p$ We set $p=1-\exp(-1)\approx 0.632$. This distribution has skewness $(2-p)/\sqrt{(}1-p)\doteq 2.25$ and kurtosis $6+p^{2}/(1-p)\doteq 7.09$. It is thus a more severe discrete distribution than $\mathrm{Pois}(1)$.
Figure 7: Same as Figure 1 except that $\Pr(x_{i}=x)=p(1-p)^{x-1}$ for $p=1-\exp(-1)$.
The results for this distribution are shown in Figure 7. All of the methods with $\Pr(w_{i}=0)>0$ had infinite RMSLs for all sample sizes $2\leqslant n\leqslant 20$. For $10\leqslant n\leqslant 20$ the Student’s $t$ intervals alternate between higher and lower coverage as $n$ changes between odd and even. The lognormal intervals do this too, except that they get shorter for odd sample sizes while Student’s $t$ intervals get shorter for even sample sizes. These alternations can be partially understood in terms of the discreteness of the distribution of $(x_{1},x_{2},\dots,x_{n})$. For example when $n=10$ there is roughly $6.7\%$ probability that the sample has one $2$, two $1$s and eight 0s. It is not clear why some of the ACIs are strongly affected by this discreteness and others are not. For instance, the power and beta bootstrap $t$ methods over-cover the true mean for $7\leqslant n\leqslant 10$ and do not show alternating coverage. Similarly, the ACI coverage levels for Poisson data do not have such a sharp alternation pattern.
6 Discussion
Nonparametric ACIs are well understood in the limit as $n\to\infty$ when $F$ satisfies some conditions based on moments and also a Cramér condition. This paper looks at small $n$, at some distributions with large or even infinite kurtosis as well as distributions with atoms that do not satisfy the Cramér condition. The conclusions from the numerical investigations are:
  1. a)
    For general purpose nonparametric ACIs for the mean of a small sample, we can rule out BCa  and the Mammen and lognormal bootstrap $t$ ACIs. They all have below nominal coverage in every example considered here while other methods did better. While BCa has advantages they do not extend to ACIs for the mean based on a small sample.
  2. b)
    Student’s $t$ intervals did well on symmetric Gaussian and heavy-tailed data, but not on skewed data. These outcomes are well understood theoretically.
  3. c)
    Methods with second order accuracy for one-sided confidence intervals were not necessarily better at two-sided coverage. The BCa, Mammen, and lognormal samplers are second order accurate but had severe under-coverage. The second order accurate beta bootstrap $t$ had worse under-coverage than the power bootstrap $t$ for several of the examples.
  4. d)
    Methods with $\Pr(v_{i}>0)=1$ never give infinite length confidence intervals. When at least two of the observations are distinct, they always have $\hat{\sigma}^{*}>0$. If all observations are equal they default to the interval $\bar{x},\bar{x}]$. These methods are especially valuable when the distribution of $x_{i}$ has atoms. In those cases the methods with $\Pr(v_{i}=0)>0$ commonly had infinite RMSLs.
  5. e)
    For the lognormal distribution, which is continuous with extremely heavy tails, the best coverage came from methods that could select $v_{i}=0$. Of those, the double-or-nothing bootstrap $t$ had the highest coverage, even over-covering for $n\leqslant 19$.
  6. f)
    The value $n=10$ is a convenient dividing line between small and very small sample sizes. For $n\leqslant 10$ the RMSL versus coverage curves frequently looked very different from any expected asymtotic behavior. For $n\geqslant 10$, most curves already looked to be following a trajectory towards the reference point.
Among methods with good asymptotic performance we should prefer those that do well on small sample sizes because that is where there can be important differences. Of the methods in this paper the power law bootstrap $t$ and the beta bootstrap $t$ consistently showed intermediate length and coverage, coming between methods that severely undercovered or had severely long intervals. Of those two, only the beta bootstrap $t$ is second order accurate.
If one has an extreme preference for coverage and much less concern for interval length, then the multinomial bootstrap $t$, double-or-nothing bootstrap, Poisson bootstrap and half sampling bootstrap provide it. Of those only the Poisson and multinomial bootstrap $t$ samplers are second order accurate.
We compared BCa, Student’s $t$ and several bootstrap $t$ algorithms. The BCa algorithm was investigated only for multinomial weights. It might benefit from some other weight distribution. However, the ACIs from BCa are nested inside the convex hull $x_{(1)},x_{(n)}]$ of the data and that already limits their coverage for small $n$. Empirical likelihood ACIs (Owen,, 2001) benefit from bootstrap calibration but are also nested inside the convex hull, though there has been some work on versions that extend beyond the convex hull of the data (Chen et al.,, 2008; Emerson and Owen,, 2009; Tsao and Wu,, 2013).
Acknowledgments
This work was supported by the National Science Foundation under grant DMS-2152780. I thank Tilmman Gneiting, Enno Mammen, Lisa Hofer and Tim Hesterberg for helpful discussions.
Appendix A Proofs
A.1 Technical lemmas
Lemma 1.
Let $v_{1},\dots,v_{n}$ be independent with $\mathbb{E}(v_{i})$ uniformly bounded away from 0 and from infinity. Suppose also that $\mathbb{E}(v_{i}^{-\alpha})\leqslant M_{\alpha}<\infty$ holds for all $i=1,\dots,n$ and some $\alpha>0$. Let $p<0$ satisfy $n\geqslant-p/\alpha$. Then
$\mathbb{E}\Bigl{(}\Bigl{(}\frac{\sum_{i=1}^{n}v_{i}}{\sum_{i=1}^{n}\mathbb{E}(v_{i})}\Bigr{)}^{p}\Bigr{)}\leqslant\Bigl{(}\frac{1}{\sqrt[\alpha]{M_{\alpha}}\min_{1\leqslant i\leqslant n}\mathbb{E}(v_{i})}\Bigr{)}^{p}.$
Proof.
This is Proposition A.3 of Hoyt and Owen, (2024). ∎
Our setting has $\mathbb{E}(v_{i})=1$ and so for $\bar{v}=(1/n)\sum_{i=1}^{n}v_{i}$, we get $\mathbb{E}\bigl{(}\bar{v}^{p}\bigr{)}\leqslant M_{\alpha}^{-p/\alpha}$ when $n\geqslant|p|/\alpha$. For the beta bootstrap $t$ of Section 2.5, we can use any $\alpha\in(0,1/2)$.
Lemma 2.
For $n\geqslant 1$, let $v_{1},\dots,v_{n}$ be IID random variables with $\mathbb{E}(v_{i}^{k})=\mu_{k}$ for $k\geqslant 1$ with $\mu_{1}=1$ and $\mu_{2}=2$. Then
$\mathbb{E}(\overline{v^{k}}(\bar{v}-1))$ $=\frac{\mu_{k+1}-\mu_{k}}{n},\quad\text{and}$ (22)
$\mathbb{E}(\overline{v^{k}}(\bar{v}-1)^{2})$ $=\frac{\mu_{k}}{n}+O\Bigl{(}\frac{1}{n^{2}}\Bigr{)}.$ (23)
Proof.
First
$\mathbb{E}(\overline{v^{k}}\bar{v})=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbb{E}(v_{i}^{k}v_{j})=\mu_{k}\mu_{1}+\frac{\mu_{k+1}-\mu_{k}\mu_{1}}{n}=\mu_{k}+\frac{\mu_{k+1}-\mu_{k}}{n}$
establishing (22). Second
$\mathbb{E}(\overline{v^{k}}\bar{v}^{2})$ $=\frac{1}{n^{3}}\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{\ell=1}^{n}\mathbb{E}(v_{i}^{k}v_{j}v_{\ell})$
$=\frac{1}{n^{3}}\bigl{(}n\mu_{k+2}+n(n-1)(\mu_{k}\mu_{2}+2\mu_{k+1}\mu_{1})+n(n-1)(n-2)\mu_{k}\mu_{1}^{2}\bigr{)}$
$=\mu_{k}+\frac{2\mu_{k+1}-\mu_{k}}{n}+\frac{\mu_{k+2}-2\mu_{k+1}+\mu_{k}}{n^{2}}$
after rearrangement. Now
$\mathbb{E}(\overline{v^{k}}(\bar{v}-1)^{2})$ $=\mu_{k}+\frac{2\mu_{k+1}-\mu_{k}}{n}+\frac{\mu_{k+2}-2\mu_{k+1}+\mu_{k}}{n^{2}}-2\Bigl{(}\mu_{k}+\frac{\mu_{k+1}-\mu_{k}}{n}\Bigr{)}+\mu_{k}$
$=\frac{\mu_{k}}{n}+O\Bigl{(}\frac{1}{n^{2}}\Bigr{)},$
establishing (23). ∎
A.2 Proof of Theorem 1
Recall that $W_{i}=nv_{i}/\sum_{i^{\prime}=1}^{n}v_{i^{\prime}}$. We begin the case of $\mathbb{E}(W_{i}^{2})$ by using a Taylor expansion of $1/\bar{v}^{2}$ around $\bar{v}=1$ getting
$\frac{\overline{v^{2}}}{\bar{v}^{2}}=\overline{v^{2}}\bigl{(}1-2(\bar{v}-1)+3(\bar{v}-1)^{2}-4(\bar{v}-1)^{3}z^{-5}\bigr{)}$ (24)
where $z$ is between $1$ and $\bar{v}$. The expected sum of the first three terms using $k=2$ in Lemma 2 is
$\mu_{2}-2\frac{\mu_{3}-\mu_{2}}{n}+3\frac{\mu_{2}}{n}+O\Bigl{(}\frac{1}{n^{2}}\Bigr{)}=\mu_{2}+\frac{10-2\mu_{2}}{n}+O\Bigl{(}\frac{1}{n^{2}}\Bigr{)}.$
It remains to bound the error term.
For $z$ between $\bar{v}$ and $1$, $1/z\leqslant 1+1/\bar{v}$. Then $\mathbb{E}(z^{-t})=O(\mathbb{E}(\bar{v}^{-t}))=O(1)$ for any $t>0$ and
$\mathbb{E}\bigl{(}\overline{v^{2}}(\bar{v}-1)^{4}z^{-5}\bigr{)}\leqslant\mathbb{E}\Bigl{(}{\overline{v^{2}}}^{\,a}\Bigr{)}^{1/a}\mathbb{E}\Bigl{(}(\bar{v}-1)^{4b}\Bigr{)}^{1/b}\mathbb{E}\Bigl{(}z^{-5c}\Bigr{)}^{1/c}$
for positive $a,b,c$ with $1/a+1/b+1/c=1$. Taking $a=2+\epsilon$, $b=2$ and $c=2+4/\epsilon$ we find that the remainder is
$O(1)O(n^{-2})\mathbb{E}(z^{-5c})^{1/c}=O(1)O(n^{-2})O(\mathbb{E}(\bar{v}^{-5c})^{1/c})=O(n^{-2})$
as $n\to\infty$, using Lemma 1 on the third factor, establishing equation (9).
For $k=3$, the Taylor expansion is
$\frac{\overline{v^{3}}}{\bar{v}^{3}}$ $=\overline{v^{3}}\bigl{(}1-3(\bar{v}-1)+6(\bar{v}-1)^{2}-10(\bar{v}-1)^{3}z^{-6}\bigr{)}$
where $z$ is between $1$ and $\bar{v}$. The expected sum of the first three terms using $k=3$ in Lemma 2 is
$\mu_{3}-3\frac{\mu_{4}-\mu_{3}}{n}+6\frac{\mu_{3}}{n}+O\Bigl{(}\frac{1}{n^{2}}\Bigr{)}=\mu_{3}+\frac{9\mu_{3}-3\mu_{4}}{n}+O\Bigl{(}\frac{1}{n^{2}}\Bigr{)}.$
The $r=3$ remainder term is at most $|\mathbb{E}\bigl{(}10\overline{v^{3}}(\bar{v}-1)^{3}z^{-6}\bigr{)}|$. We can show that it is $O(n^{-3/2})$ similarly to the prior case, this time using $a=3+\epsilon$, $b=2$ and $c=(3+\epsilon)/(2+\epsilon)$ to establish (10).
A.3 Proof of Theorem 2
Let $S\subseteq\{1,2,\dots,n\}$ have $k\geqslant 3$ distinct values of $x_{i}$ and let $\delta$ be the smallest value of $|x_{i}-x_{j}|$ for distinct $i,j\in S$. Now
$(t^{*})^{2}$ $=\frac{n\sum_{i=1}^{n}\sum_{j=1}^{n}v_{i}v_{j}(x_{i}-\bar{x})(x_{j}-\bar{x})}{\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{n}v_{i}v_{j}(x_{i}-x_{j})^{2}},$
where we can take $v_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Beta}(1/2,3/2)$ as the factor of 4 cancels between numerator and denominator. The numerator is bounded, and by the AM-GM inequality the denominator is no smaller than
$\frac{\delta^{2}}{2}\sum_{i\in S}\sum_{j\in S\setminus\{i\}}v_{i}v_{j}\leqslant\frac{k(k-1)\delta^{2}}{2}\prod_{i\in S}\prod_{j\in S\setminus\{i\}}(v_{i}v_{j})^{1/k(k-1)}=\frac{k(k-1)\delta^{2}}{2}\prod_{i\in S}v_{i}^{1/k}.$
Letting $C$ be the upper bound for the numerator
$\mathbb{E}((t^{*})^{2})\leqslant\frac{2C}{k(k-1)\delta^{2}}\mathbb{E}(v_{1}^{-1/k})^{k}.$
Now $\mathbb{E}(v_{1}^{-1/k})=\int_{0}^{1}v^{-1/2-1/k}(1-v)^{1/2}/B(1/2,3/2)\,\mathrm{d}v<\infty$ for $k\geqslant 3$. The power bootstrap $t$ replaces $\mathrm{Beta}(1/2,3/2)$ by $\mathrm{Beta}(a,1)$ for $a=\sqrt{2}-1$. It then suffices to have a11/k>1a-1-1/k>-1, which also holds for $k\geqslant 3$.
A.4 Proof of Theorem 3
Relabel the observations if necessary so that $x_{i}=x_{1}$ for $1\leqslant i\leqslant n_{1}$ and $x_{i}=x_{n}$ for $n_{1}<i\leqslant n$. Letting $\lambda=n_{1}/n$ and $\lambda^{*}=\sum_{i=1}^{n_{1}}w_{i}$, we find that $\bar{x}^{*}-\bar{x}=(\lambda^{*}-\lambda)(x_{1}-x_{n})$ and $\hat{\sigma}^{*2}=\lambda^{*}(1-\lambda^{*})(x_{1}-x_{n})^{2}$. Then
$t^{*}=\sqrt{n}\frac{\lambda^{*}-\lambda}{\sqrt{\lambda^{*}(1-\lambda^{*})}}\mathrm{sign}(x_{1}-x_{n}).$ (25)
For $j\in\{1,n\}$ let $u_{j}=\sum_{i=1}^{n}v_{i}1\{x_{i}=x_{j}\}$ for $v_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Beta}(1/2,3/2)$, so
$(t^{*})^{2}$ $=n\frac{(\lambda^{*}-\lambda)^{2}}{\lambda^{*}(1-\lambda^{*})}\leqslant\frac{n}{\lambda^{*}(1-\lambda^{*})}=\frac{n(u_{1}+u_{n})^{2}}{u_{1}u_{n}}=n\Bigl{(}\frac{1}{u_{1}}+\frac{1}{u_{n}}+2\Bigr{)}.$
Finally, $\mathbb{E}(1/u_{1})<\infty$ if $\int_{0}^{1}v^{-1/2-1/n_{1}}(1-v)^{1/2}\,\mathrm{d}v<\infty$, so $\mathbb{E}((t^{*})^{2}\!\mid\!x_{1},\dots,x_{n})<\infty$ holds if $\min(n_{1},n-n_{1})\geqslant 3$.
Appendix B Literature on small $n$ bootstrap
While most of the theory for the bootstrap is asymptotic as $n\to\infty$ there have also been some combinatorial investigations of the bootstrap that pertain to small $n$.
Fisher and Hall, (1991) investigate many small sample issues. Suppose that all $x_{i}$ are distinct as must happen when $F$ has no atoms. They note that $\Pr(\hat{\sigma}^{*}=0)=n^{1-n}$. Also there are $2n-1\choose n$} distinct multinomial count vectors $\boldsymbol{v}=(v_{1},\dots,v_{n})$. The most likely of these is $(1,1,\dots,1)$ with probability $n!/n^{n}$. The least likely count vectors have only one nonzero element. There are $n$ such outcomes with probability $n^{-n}$ each.
Let the sorted values of $v_{i}$ be $v_{[1]}\geqslant v_{[2]}\geqslant\cdots\geqslant v_{[1]}$. Fisher and Hall, (1991) take a special interest in outcomes where $v_{[1]}$ is n1n-1 or n2n-2 as those can give very small values of $\hat{\sigma}^{*}$. They find that
$\Pr(v_{[1]}=n-1)$ $=\Bigl{(}1-\frac{1}{n}\Bigr{)}n^{3-n}$ for $n\geqslant 3$
$\Pr(v_{[1]}=n-2,v_{[2]}=v_{[3]}=1)$ $=\frac{1}{2}\Bigl{(}1-\frac{1}{n}\Bigr{)}^{2}\Bigl{(}1-\frac{2}{n}\Bigr{)}n^{5-n}$ for $n\geqslant 4$
$\Pr(v_{[1]}=n-2,v_{[2]}=2)$ $=\frac{1}{2}\Bigl{(}1-\frac{1}{n}\Bigr{)}^{2}n^{4-n}$ $\text{for $n\geqslant 5$}.$
They tabulate the probability that some value appears at least n2n-2 out of $n$ times for $n\geqslant 5$. They find that it requires $n\geqslant 7$ for this probability to be below $0.05$. They then say “Therefore we may expect to experience problems with tied resample values when using the percentile-$t$ method in samples of size $n\leqslant 6$.” Those tied resample values do serve to make the bootstrap $t$ intervals wider than they otherwise would be and hence increase their coverage.
Fisher and Hall, (1991) adopt a convention for $\hat{\sigma}^{*}=0$. Translating to the present notation, their equation (2.2) interprets $t^{*}=(\bar{x}^{*}-\bar{x})/\hat{\sigma}^{*}\leqslant Q$ as
$\bar{x}^{*}-\bar{x}\leqslant\hat{\sigma}^{*}Q.$ (26)
Now if $\hat{\sigma}^{*}=0$ and $\bar{x}^{*}<\bar{x}$ then (26) holds for every real $Q$ no matter how small, so we should take $t^{*}=$-$\infty$. Similarly for $\bar{x}^{*}>\bar{x}$, equation (26) is false no matter how large we make $Q$ and so we should take $t^{*}=\infty$. This matches our convention of taking $t^{*}$ to be $\infty$ when the numerator is positive and the denominator is zero and -$\infty$ when we divide a negative number by 0.
They do not mention a convention for the case where $\bar{x}^{*}-\bar{x}=\hat{\sigma}^{*}=0$. That can only happen when there are ties in the data and they assumed continuously distributed data. They describe ties as arising from rounding which they can then undo by adding a uniform variable to each $x^{*}_{i}$ in each bootstrap sample. Count data are discrete and not well described by rounding of continuous data.
Diaconis and Holmes, (1994) show how to use Gray codes to efficiently enumerate all of the bootstrap sampling counts $\boldsymbol{v}=(v_{1},\dots,v_{n})$. This allows an exact bootstrap calculation to be done by assigning the appropriate weight to the statistic computed from each of the $2n-1\choose n$} different count vectors, instead of resampling $B$ times. In the Gray code ordering any two consecutive count vectors differ in only two places allowing some efficient updates to be used.
Appendix C Conventions and edge cases
There are some odd cases that arise where we need to define a convention. These are nuisances that can arise for small $n$ but have exponentially small probability of happening as $n$ grows. We need to handle them when working with small $n$.
First, when $\hat{\sigma}^{*}=0$ and $\bar{x}^{*}\neq\bar{x}$ then we get $t^{*}=\pm\infty$. Our ACIs will still have a finite lower confidence limit if $t^{*}=\infty$ happens fewer than $B\alpha_{1}$ times and similarly for the upper limits. For some methods the $B\to\infty$ version of the bootstrap $t$ interval truly does have infinite length. We choose to accept that as an (undesirable) property of any such method instead of tweaking the method to avoid that problem. If the true length is infinite for some value of $n$, that shows that the method is not suitable for such $n$.
Next, it is possible to have $\bar{x}^{*}=\bar{x}$ at the same time that $\hat{\sigma}^{*}=0$. This can happen when one of the $x_{i}=\bar{x}$. This has positive probability for count data. Then $t^{*}=0/0$ and we must choose what to do. One choice is to deliver $t^{*}=0$. A second choice is to split the resampled $t^{*}$ into two $t^{*}$’s one equal to each of $\pm\infty$ getting weight $1/(2B)$ each when forming quantiles. This is too cumbersome and not clearly better, so when $\bar{x}^{*}=\bar{x}$ and $\hat{\sigma}^{*}=0$ we take $t^{*}=0$ by convention.
When $\rho=\Pr(v=0)>0$ and the $v_{i}$ are IID, we can get $v_{1}=v_{2}=\cdots=v_{n}=0$. Then there is no good way to define the weights $w_{i}$. Our convention is to discard such weight vectors and keep sampling until we have $B$ pseudo-count vectors that all have $\sum_{i=1}^{n}v_{i}>0$. The resulting $w_{i}$ are still exchangeable. An alternative is to take $t^{*}=0$ for such cases, as if $w_{i}$ were all equal, and hence equal to $1/n$. The probability that this event happens is $\rho^{n}$, which equals $\exp(-n)$ for the Poisson bootstrap and $2^{-n}$ for double-or-nothing.
In multinomial sampling, there is probability $n^{1-n}$ that all sample indices $j(i)$ are identical. In that case we will get $\hat{\sigma}^{*}=0$ even when $x_{i}$ are all distinct. If $\hat{\sigma}^{*}=0$ often enough, then the bootstrap $t$ ACI can have infinite length. We can ensure that the true length of the bootstrap $t$ ACI is finite, for $x_{i}$ with a continuous distribution, by taking $n^{1-n}<\min(\alpha_{1},1-\alpha_{2})$. For $\alpha_{1}=1-\alpha_{2}=0.025$ we require $n\geqslant 4$. When $n=4$ and $B=2000$, we need $B\times 0.025=50$ infinite $t^{*}$ values to get an infinite sample quantile. Then $\Pr(\mathrm{Bin}(2000,4^{-3})\geqslant 50)\doteq 0.0011$ so $\hat{\sigma}^{*}=0$ is moderately rare for $n=4$ and since $\Pr(\mathrm{Bin}(2000,5^{-4})\geqslant 50)\doteq 1.5\times 10^{-41}$ is a non-issue for continously distributed data with $n\geqslant 5$. From here on, we only consider the $B\to\infty$ case.
The above analysis is somewhat conservative because continuous data will have $x_{(1)}<\bar{x}<x_{(n)}$ with probability one. Then the events with $\hat{\sigma}^{*}=0$ will be split between some that give $t^{*}=$-$\infty$ and others that give $t^{*}=+\infty$ and an infinite interval only arises when one of those sub-populations has probability $0.025$ or more. When $x_{(j)}<\bar{x}<x_{(j+1)}$, then an infinite lower limit has probability $jn^{-n}$ in multinomial sampling, and so $\Pr(t^{*}=-\infty\!\mid\!x_{1},\dots,x_{n})\leqslant(n-1)/n^{n}$ which is below $0.025$ for $n\geqslant 4$. By symmetry $\Pr(t^{*}=\infty\!\mid\!x_{1},\dots,x_{n})<0.025$ also holds for $n\geqslant 4$.
For the double-or-nothing bootstrap, $\Pr(Y=1)=n2^{-n}$ and $\Pr(Y=1\!\mid\!Y>0)=n2^{-n}/(1-2^{-n})$. Both of these become below $0.025$ once $n\geqslant 9$. The same holds for $(n-1)2^{-n}$ and $(n-1)2^{-n}/(1-2^{-n})$.
For the Poisson bootstrap $\Pr(Y=1)=n\sum_{x=0}^{\infty}(e^{-1}/x!)^{n}$. With $n\geqslant 7$ and $B=\infty$, $\Pr(|t^{*}|=\infty\!\mid\!x_{1},\dots,x_{n})=0$ when $x_{i}$ are distinct. If we don’t discard cases with $Y=0$ then this holds for $n\geqslant 6$.
The BCa method can never give us an ACI of infinite length because those intervals are always subsets of $x_{(1)},x_{(n)}]$. The BCa intervals are not well defined when $x_{1}=x_{2}=\cdots=x_{n}$. In that case we take the ACI to be $\bar{x},\bar{x}]$. One would not normally seek a confidence interval for the mean of a sample where all the values were equal. However in a high throughput setting it is worthwhile to have a default way to handle such irregular cases.